### v2

# Hierarchical Clustering Algorithms on Poisson and Cox Point Processes

Abstract
Clustering is a widely used technique in unsupervised learning to identify groups within a dataset based on the similarities between its elements. This paper introduces three new hierarchical clustering models, Clustroid Hierarchical Nearest Neighbor ($\mathrm{CHN}^{2}$), Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$), and Hausdorff (Complete Linkage) Hierarchical Nearest Neighbor ($\mathrm{H}^{2}\mathrm{N}^{2}$), all designed for datasets with a countably infinite number of points. These algorithms proceed through multiple levels of clustering and construct clusters by connecting nearest-neighbor points or clusters, but differ in the distance metrics they employ (clustroid, single linkage, or Hausdorff, respectively). Each method is first applied to the homogeneous Poisson point process on the Euclidean space, where it defines a phylogenetic forest, which is a factor of the point process and therefore unimodular. The results established for the $\mathrm{CHN}^{2}$ algorithm include the almost-sure finiteness of the clusters and bounds on the mean cluster size at each level of the algorithm. The mean size of the typical cluster is shown to be infinite. Moreover, the limiting structure of all three algorithms is examined as the number of levels tends to infinity, and properties such as the one-endedness of the limiting connected components are derived. In the specific case of $\mathrm{SHN}^{2}$ on the Poisson point process, the limiting graph is shown to be a subgraph of the Minimal Spanning Forest. The $\mathrm{CHN}^{2}$ algorithm is also extended beyond the Poisson setting, to certain stationary Cox point processes and and more generally to all stationary point process with no second-order descending chains. Similar finite-cluster properties are shown to hold in these cases. It is also shown that efficient detection of Cox-triggered aggregation can be achieved through this clustering algorithm.
1 Introduction
Clustering is a popular method in unsupervised learning, which plays an important role in data science. Given a representation of the objects, clustering aims at identifying groups, or clusters, based on some similarity measure between them. Usually, objects are represented as a finite number of points in $\mathbb{R}^{d}$.
Clustering algorithms can be organized into two broad categories: point assignment and hierarchical. In point assignment algorithms, some clusters are predefined, points are considered in some order, and each one is assigned to the cluster into which it best fits. Here, an important parameter is the number of clusters. The best known point assignment algorithm is the $k$-means algorithm (see [26] for a definition and further details). Hierarchical algorithms start with each point forming its own cluster. Clusters are then connected/merged on the basis of their similarities and form new clusters. The process of connecting and merging clusters is repeated until one reaches a single cluster to which all the points belong. A key component of hierarchical clustering is the function used to measure distances between clusters. This includes minimum distance (single linkage), Hausdorff distance (complete linkage), and centroid/clustroid distances. In centroid/clustroid distances, one or more representative points within each cluster, referred to as centroid or clustroid [^1] When the representative point of a cluster is not precisely located at the center of the cluster’s data, the term clustroid is used instead of centroid., are selected to collectively represent the entire cluster. The distance between clusters is then calculated on the basis of the distances between their respective centroids (clustroids). For more details, see [21] and [15].
The aim of the present paper is to study the properties of a class of natural clustering algorithms that essentially belong to the hierarchical category but also share features of point assignment. This class comprises the three hierarchical clustering algorithms listed in the abstract, namely the $\mathrm{CHN}^{2}$, $\mathrm{SHN}^{2}$, and $\mathrm{H}^{2}\mathrm{N}^{2}$ clustering algorithms. These algorithms follow the same overall principles for constructing clusters at each level, but differ in the distance used between clusters: $\mathrm{CHN}^{2}$ relies on clustroid distance, $\mathrm{SHN}^{2}$ uses single linkage distance, whereas $\mathrm{H}^{2}\mathrm{N}^{2}$ employs the Hausdorff (complete linkage) distance. The main focus of this paper is on the properties of $\mathrm{CHN}^{2}$. Some properties of $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ are however also discussed in Section 3.
In what follows, we give a high-level description of the clustering algorithms, centered on the $\mathrm{CHN}^{2}$ case, where the clustroid of a cluster is chosen as the pair of points that achieves the minimum distance in the cluster. More details on the definition of $\mathrm{CHN}^{2}$ are provided in Section 2. The variants associated with $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ are described in Section 3. Below, $D$ is the dataset, consisting of a countable collection of points in $\mathbb{R}^{d}$. The precise underlying assumptions will be spelled out later. They will for instance be satisfied by the collection of atoms of a homogeneous Poisson point process of the Euclidean space.
At level zero, each point in $D$ connects to its nearest neighbor point. Under these assumptions, this defines a collection of finite connected components called clusters of order 0 on the data set. Each cluster is made up of two directed trees, called cluster subtrees of order 0, that are connected to a directed cycle containing a pair of mutually nearest neighbor points (MNN). Call each such pair a cycle of order 0 and the points belonging to a cycle of order 0 cluster heads of order 0. See Figure 1  [^2] The graph constructed at this level is classical and called the nearest neighbor graph on $D$. Nearest neighbor graphs have been extensively studied in computational geometry ([12], and [13]) and play an important role in clustering and classification methods (see [9], [21], and [16]). They are known to capturing local proximity structures effectively, which makes them valuable in manifold learning, spatial data analysis, and other tasks where local neighborhoods are of interest (see [7])..
Figure 1: Clusters of order zero generated by the $\mathrm{CHN}^{2}$ clustering algorithm on a finite dataset. The blue arrows connect each point to its nearest neighbor. The red cycles represent cycles of order zero, connecting points that are mutual nearest neighbors. The details of the elements within one cluster are shown in the figure.
At level one each 0-cycle is connected to its nearest 0-cycle. The pseudo-distance between two cycles is defined as the minimum (or infimum) distance between any point in the first 0-cycle and any point in the second. In the literature on clustering algorithms, this pseudo-distance is called the single linkage distance. The last name will be used in this paper to indicate this pseudo-distance. More precisely, define the exit point of a 0-cycle as the point in this 0-cycle that is closest to the nearest 0-cycle. For each exit point, replace the initial directed connection to the other point in its own cycle by a directed connection to the closest point in the nearest 0-cycle. This creates a directed graph on cluster heads of order 0, with finite connected components. Each component in this graph consists in directed trees whose nodes are cluster heads of order 0. Each tree is connected to a cycle containing a pair of 0-cluster heads that belong to two 0-cycles which are mutual nearest neighbors for single linkage. Each connected component of this graph, along with the cluster subtrees of order zero attached to them, defines a cluster of order 1 on the dataset. Each order 1 cluster contains some directed trees, called cluster subtrees of order 1. The nodes of these trees are either non-MNN points or cluster heads of order 0. Each such tree is connected to a cycle of length 2, which connects two cluster heads of order 0 which belong to two mutual nearest neighbor 0-cycles. This pair forms a cycle of order 1 and the points belonging to such a cycle will be called cluster heads of order 1. The latter are hence a subclass of that of cluster heads of order 0. See Figure 2.
Figure 2: Clusters of order one generated by the $\mathrm{CHN}^{2}$ clustering algorithm on a finite dataset. Each cycle of order zero is connected to its nearest neighbor pair with respect to the single linkage distance. The green arrows indicate the new connections of the vertices that are exit points of a cycle of order zero. The pink points represent the cluster heads of order 1, which are the exit points of the mutual nearest neighbor cycles with respect to the single linkage distance. The details of the elements within one cluster are shown in the figure.
By iterating this procedure, one can proceed hierarchically as above to define exit points of order 1 or higher, clusters of order 2 or higher, cluster subtrees of order 2 or higher, as well as cycles and cluster heads of order 2 or higher. For all $k\geq 0$, this construction results in a directed graph of order $k$ where each component or cluster of order $k$ is finite, made of a collection of cluster subtrees of order $k$, connected to a $k$-cycle connecting two cluster heads of order $k$. A cluster head of order $k$ is an exit point of order $k-1$ that connects two mutual nearest neighbor cycles of order $k-1$ with respect to the single linkage distance. Cluster heads of order $k$ hence form a subclass of those of order $k-1$.
All classification methods receive a phylogenetic interpretation, where nodes are seen as species. In the $\mathrm{CHN}^{2}$ clustering algorithm, the directed graph of order $k$ organizes species into an infinite collection of finite $k$-clusters. Each cluster contains directed trees (subtrees of order $k$) that specify parent-child relationships among species, up to cluster heads of order $k$, which can be considered the archetypes of the cluster. For these cluster heads, it is not possible to determine the parent species without increasing the order $k$.
Here are a few observations on $\mathrm{CHN}^{2}$ clustering:
  • It is defined as a hierarchical algorithm; however, it can also be considered a point assignment algorithm because the cluster heads of a given order are, in a sense, predefined through a mechanism that does not require the whole dataset but only the cycles of the previous order and the computation of the mutual nearest neighbor pairs of cycles.
  • For all $k$, the cluster heads (which can be seen as archetypes) of level $k$ are a deterministic functional of the data set. Their definition involves no extra randomness, as in, for example, the techniques based on the thinning of point processes described in more detail in Section 6.
  • $\mathrm{CHN}^{2}$ clustering is applicable to infinite data sets. Additionally, using the pseudo-distance between pairs results in limited computational effort, making it well-suited for large datasets. This feature distinguishes it from other known distance-based hierarchical clustering algorithms, such as the nearest-neighbor chain (see [8] and [18]).
  • Furthermore, the $\mathrm{CHN}^{2}$ clustering algorithm supports parallelization, allowing trees to be built in two sufficiently separated regions simultaneously. This enables faster processing and improved scalability.
As mentioned above, the aim of this paper goes beyond introducing new hierarchical clustering algorithms. It also studies how the algorithms behave when applied to point processes. Here is a summary of the main analytical results:
First, in Section 2, the $\mathrm{CHN}^{2}$ algorithm is considered on a homogeneous Poisson point process (PPP). It is shown that the algorithm can be described as a sequence of deterministic point-shifts on the PPP, with each point-shift representing a step of the algorithm (for more details on point-shifts, see [5]). At each step $k$, the point-shift constructs a random directed graph on the PPP. One key property is that the connected components (or clusters) of the directed graph of each level $k\in\mathbb{N}$ of the algorithm are all a.s. finite, as shown in Theorem 2.9. This theorem is proven in Section 5.1 by introducing the notion of Second-Order Descending Chain, which generalizes that of a descending chain in point processes (see [11] and [22]). In this chain, the distance between two consecutive points $x_{k}$ and $x_{k+1}$, denoted as $d_{k}$, satisfies the relation $d_{k}\leq\max(d_{k-1},d_{k-2})$, for all $k\geq 3$. It is shown that the PPP does not admit an infinite second-order descending chain a.s. This is shown to imply that every level of the $\mathrm{CHN}^{2}$ clustering algorithm contains only finite clusters. So the algorithm remains well-defined as it transitions from level $k$ to level $k+1$. Moreover, by considering the exit points at level $k$ as representatives of each cluster, it is shown that the intensity of the exit points at level $k$ is at most $\rho/2^{k+1}$, where $\rho$ is the intensity of the cluster heads of level 0 (which is known in closed form). Simulation results indicate that this upper bound can be further refined to approximately $\rho/3^{k+1}$ (see Figure 5). This result provides insight into the speed at which clusters merge at each step of the algorithm, forming new clusters at subsequent levels, see Remark 2.6.
Moreover, at each level, the resulting random graph is unimodular (see [2]). Corollary 2.8 further shows that, as $k\to\infty$, the limiting point-shift called Clustroid Hierarchical Nearest Neighbor $(\mathrm{CHN}^{2})$ Point-Shift exists. This limiting (point-shift) graph is called the Clustroid Hierarchical Nearest Neighbor Eternal Family Forest ($\mathrm{CHN}^{2}$ EFF). This infinite graph is unimodular under the Palm probability of $\Phi^{0}$.
The $\mathrm{CHN}^{2}$ EFF forms an infinite spanning forest or tree on the PPP. All connected components of $\mathrm{CHN}^{2}$ EFF have the same distribution (Proposition 2.11), and Corollary 2.12 shows that each connected component in $\mathrm{CHN}^{2}$ EFF is one-ended. These results hold for any stationary point process that does not admit a second-order descending chain almost surely (see Sections 2.2 and 5.1 for further details).
As mentioned above alternative definitions of the distance between clusters are also possible. Specifically, in Section 3, the two variants of the algorithm alluded to above are introduced more precisely: the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) and the Hausdorff (Complete Linkage) Hierarchical Nearest Neighbor ($\mathrm{H}^{2}\mathrm{N}^{2}$) algorithms, which employ single linkage and Hausdorff distances, respectively. The properties of these algorithms on Poisson point processes can be analyzed by techniques similar to those of the $\mathrm{CHN}^{2}$ clustering algorithm.
In these cases, it is necessary to generalize the definition of the algorithm at each level to handle potentially infinite clusters. The random graphs constructed at each level of the $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithms are unimodular, and their weak limits exist. These limits are referred to as the $\mathrm{SHN}^{2}$ Spanning Forest and the $\mathrm{H}^{2}\mathrm{N}^{2}$ Spanning Forest, respectively (see Theorem 3.2). To prove the existence of these forests, a sequence of point-shifts on the corresponding graphs is defined in Section 5.1. It is shown that the weak limits of these point-shifts exist and that the associated graphs are one-ended (see Remark 5.12). Moreover, Corollary 3.4 shows that, on the PPP, the $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the Minimal Spanning Forest (MSF) introduced by Aldous and Steele in [23]. This result leads to a new class of open problems on the $\mathrm{SHN}^{2}$ Spanning Forest, potentially offering further insights into the behavior of the MSF on the PPP (see Section 7).
One of the generic goals of all clustering algorithms is to detect aggregation of points in a dataset. Working with a PPP is useful in this context because the PPP can be viewed as a null hypothesis, i.e., a baseline model that assumes no underlying aggregation other than those triggered by the fluctuations of the homogeneous PPP. To examine the behavior of the $\mathrm{CHN}^{2}$ algorithm in scenarios where aggregation does occur, Section 4 applies this algorithm to instances of Cox point processes, which exhibit noticeable aggregation (see Figures 11 and 12). It is shown that, in these cases, the $\mathrm{CHN}^{2}$ algorithm enjoys properties analogous to those observed under the PPP, such as the finiteness of clusters at each hierarchical level and the weak convergence of the limiting graph. Furthermore, it is shown how the $\mathrm{CHN}^{2}$ clustering algorithm effectively detects aggregation in the dataset.
2 Clustroid Hierarchical Nearest Neighbor Clustering on a Point Process
This section considers the $\mathrm{CHN}^{2}$ algorithm on the homogeneous PPP. As previously stated in the introduction, each level of the algorithm can be characterized as a point-shift on the PPP. A point-shift $f$ maps, in a translation invariant way, each point of a stationary point process $\phi$ to some point of $\phi$ (see [20], [24], and [4] for more details).
The homogeneous PPP in $\mathbb{R}^{d}$ is denoted by
$\Phi^{0}=\sum_{i}\delta_{x_{i}}.$
The construction described below is scale-invariant, so there is no restriction in assuming that its intensity is $1$.
2.1 Construction of the Pre-limit Point-Shifts
We construct a sequence $f^{n}$, $n\geq 0$, of point-shifts on $\Phi^{0}$ by induction. For this, we will use the following pseudo-distance function acting on two sets of points in $\mathbb{R}^{d}$:
$\delta(S,T)=\min_{x\in S,y\in T}d(x,y),$ (1)
where $d$ denotes the Euclidean distance. As mentioned in the introduction, this pseudo-distance is called the single linkage distance in the literature. This function is non-negative and symmetric. However, it does not satisfy the triangle inequality. Additionally, different sets have a pseudo-distance of 0 when they share at least one point.
Order 0
The first point-shift, $f^{0}$, is the nearest neighbor point-shift on $\Phi^{0}$ w.r.t. Euclidean distance $d$ between points. That is for all $x$, $f^{0}(x)={\mathrm{NN0}}(x)$, where NN0 maps a point to its nearest neighbor in $\Phi^{0}$.
Proposition 2.1.
The graph of the point-shift $f^{0}$ (the $f^{0}$-graph) has all its connected components in the $\mathcal{F/F}$ class in the sense of the point-shift classification theorem in [6]. That is, all of its connected components are a.s. finite and have a unique cycle.
Proof.
Note that any infinite path in the $f^{0}$-graph on the PPP is a descending chain. A descending chain in a point process is an infinite sequence of points in the process where the distances between consecutive points form a decreasing sequence (see [22], [17], and [11]). It is shown in [11] that the homogeneous PPP does not admit any descending chains a.s. Therefore, the connected components of the $f^{0}$-graph are almost surely finite and belong to the $\mathcal{F/F}$ class of the point-shift classification theorem. Each connected component consists of directed trees connected to a single cycle that connects two mutual nearest neighbor (MNN) points of $\Phi^{0}$. ∎
Definition 2.2.
Each connected component of the $f^{0}$-graph on the PPP has a cycle of length two. These cycles are called 00-cycles. The connected components of the $f^{0}$-graph are also referred to as clusters of order zero. The vertices in the $f^{0}$-graph that belong to 00-cycles are called cluster heads of order zero. The union of cluster heads of order zero forms a sub-PP of $\Phi^{0}$ denoted by $\Phi^{0}_{c}$. Let $\Phi^{0}_{a}=\Phi^{0}\setminus\Phi^{0}_{c}$.
If the edges of the 00-cycles in the $f^{0}$-graph are deleted, one obtains a collection of directed trees referred to as the cluster subtrees of order 0. Each vertex $x$ in a cluster subtree of order 0 is either a cluster head of order zero (i.e., $x\in\Phi^{0}_{c}$) or is such that $x\in\Phi^{0}_{a}$, and there is a path from $x$ to a cluster head of order 0. See Figure 3.
Figure 3: Left picture: The $f^{0}$-graph generated on PPP in blue with the 00-cycles shown in bold;
Right picture: Cluster subtrees of order 0 obtained by deleting the edges of the 00-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order zero, and the other vertices have directed paths towards the cluster head of order zero within the cluster.
Order 1
Before proceeding to the definition of the point-shift $f^{1}$, we need the following definitions:
Definition 2.3.
Let $\{C^{0}_{i}\}_{i}$ denote the collection of 00-cycles, and let $\{S^{0}_{i}\}_{i}$ be the collection of pairs of points in these cycles, namely $S^{0}_{i}$ is the pair of head points belonging to the cycle $C^{0}_{i}$. The sets $\{S^{0}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{0}_{c}$.
Consider the nearest neighbor map, denoted by NN1, on $\{S^{0}_{i}\}_{i}$, which maps each pair $S^{0}_{i}$ to its nearest neighbor pair in $\{S^{0}_{j}\}_{j}$. The distance between the sets is considered to be the single linkage distance, $\delta$, defined in (1). Specifically, $S_{i}^{0}$ maps to $S_{j}^{0}$ if $\delta(S_{j}^{0},S_{i}^{0})<\delta(S_{k}^{0},S_{i}^{0})$ for all $k\neq i,j$.
[…]
Definition 2.4.
[…]
The vertices in the $f^{1}$-graph that belong to $1$-cycles are called cluster heads of order 1. The union of cluster heads of order 1 forms a sub-PP of $\Phi_{c}^{0}$, denoted by $\Phi^{1}_{c}$. Let $\Phi^{1}_{a}=\Phi^{0}\setminus\Phi^{1}_{c}$.
If the edges of the $1$-cycles in the $f^{1}$-graph are deleted, one obtains a collection of directed trees referred to as cluster subtrees of order 1. Each vertex $x$ in a cluster subtree of order 1 is either a cluster head of order 1 (i.e., $x\in\Phi^{1}_{c}$) or such that $x\in\Phi^{1}_{a}$, and there is a path from $x$ to a cluster head of order 1. See Figure 4.
Figure 4: Left picture: The $f^{1}$-graph generated on the PPP with the $1$-cycles shown in green bold. The difference between this graph and the $f^{0}$-graph is the addition of new edges from a cycle of order zero to the closest point of the nearest cycle of order zero, shown in green.
Right picture: Cluster subtrees of order 1 obtained by deleting the edges of the $1$-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order 1, and the other vertices have directed paths towards the cluster head of order 1 within the cluster.
Figure 5: Estimation of the (logarithm of the) intensity of the exit points at different levels of the algorithm. The initial point process contains approximately 12,000 points. As shown in the plot, the intensity decays by a factor close to $1/3$ at each level, suggesting that the intensity of the exit points at level $k$ is approximately $(1/3)^{k}$.
Order 2
To define the next point-shift $f^{2}$ on the PPP, we need to introduce some notions similar to those defined earlier in Definition 2.3.
Definition 2.5.
Let $\{C^{1}_{i}\}_{i}$ denote the collection of $1$-cycles, and let $\{S^{1}_{i}\}_{i}$ be the collection of all such pairs of points: $S^{1}_{i}$ is the pair of head points belonging to the cycle $C^{1}_{i}$. The sets $\{S^{1}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{1}_{c}$.
Consider the nearest neighbor map, denoted by NN2, on $\{S^{1}_{i}\}_{i}$, which maps each pair $S^{1}_{i}$ to its nearest neighbor in $\{S^{1}_{i}\}_{i}$. The distance between the sets is again the single linkage distance, $\delta$, defined in (1).
[…]
Define $f^{2}$ to be the point-shift on $\Phi^{0}$ that coincides with $f^{1}$ everywhere except for the exit points of order 1. For each such point, say $x\in S^{1}$, where $S^{1}$ is the pair where $x$ is the exit point, define $f^{1}(x)=\mathrm{nn2}(x)$. Again, $f^{2}$ defines a point-shift on $\Phi^{0}$. Its graph is $\mathcal{F/F}$, with cycles of length 2 a.s., as there are no ties a.s. The points belonging to these $2$-cycles are exit points of order 1 which form a sub PP of $\Phi^{1}_{c}$ that will be called cluster heads of order 2 and denoted $\Phi_{c}^{2}$. Let $\Phi_{a}^{2}:=\Phi^{0}\setminus\Phi_{c}^{2}$ denote the vertices of the acyclic part of the $f^{2}$ graph. When deleting the $2$-cycles, i.e. the two directed edges, one gets a collection of finite trees which will be called the cluster subtrees of order 2. See Figure 6.
Figure 6: Left picture: The $f^{2}$-graph generated on the PPP with the $2$-cycles shown in red bold. The difference between this graph and the $f^{1}$-graph is the addition of new edges from a cycle of order 1 to the closest point of the nearest cycle of order 1, shown in green. Right picture: Cluster subtrees of order 2 obtained by deleting the edges of the $1$-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order 2, and the other vertices have directed paths towards the cluster head of order 2 within the cluster.
As illustrated in the definitions of the point-shifts $f^{1}$ and $f^{2}$, for each $k\geq 2$, the point-shift $f^{k}$ can be defined using the previous level point-shift, $f^{k-1}$. More generally, $f^{k}$ is defined by induction as follows:
Induction Assumption
The induction assumption is that:
  • $f^{k-1}$ has been constructed as an $\mathcal{F/F}$ point-shift on $\Phi^{0}$ (in the case of PPP, this property is shown in Theorem 2.9);
  • The ${k-1}$-cycles $\{C^{k-1}_{i}\}$ of the connected components are all of length 2 a.s.;
  • The point process $\Phi^{k-1}_{c}$ of cyclic points of the $f^{k-1}$-graph has an intensity bounded above by $\rho/2^{k-1}$.
Induction step
Let $\{S^{k-1}_{i}\}_{i}$ be the collection pairs of points of $C^{k-1}_{i}$. Let NNk be the nearest neighbor map on $\{S^{k-1}_{i}\}_{i}$, which maps each pair $S^{k-1}_{i}$ to its nearest neighbor pair in $\{S^{k-1}_{i}\}_{i}$, again with respect to the single linkage distance, $\delta$. The NNk graph is the directed graph on $\{S^{k-1}_{i}\}_{i}$ with an edge from $S^{k-1}_{i}$ to $\mathrm{NNk}(S^{k-1}_{i})$, the nearest neighbor of $S^{k-1}_{i}$ based on the single linkage distance.
Let $\mathrm{NNk}(S^{k-1}_{i})=S^{k-1}_{j}$. There is a point, denoted $x_{i}$, in $S^{k-1}_{i}$ and a point, $x_{j}$, in $S^{k-1}_{j}$ that achieve the minimum in $\delta(S^{k-1}_{i},S^{k-1}_{j})$, denoted by $\mathrm{nnk}(x_{i})=x_{j}$. The point $x_{i}$ is called the exit point of order k-1 in the pair $S^{k-1}_{i}$.
One defines $f^{k}$ from $f^{k-1}$ by keeping the images of all points the same except for exit points of order $k-1$. The image of such a point say $x\in S^{k-1}_{i}$ define $f^{k}(x)=\mathrm{nnk}(x)$.
Each connected component of the graph of $f^{k}$ on $\Phi^{0}$ is referred to as a cluster of order k. It will be shown in Corollary 5.7 that these connected components are almost surely $\mathcal{F/F}$, meaning that each cluster in the graph of $f^{k}$ is finite and contains a unique cycle. These cycles are of length 2 a.s. and connect two exit points of order $k-1$, denoted $x$ and $y$, and such that $\mathrm{nnk}(x)=y$ and $\mathrm{nnk}(y)=x$. These cycles are called $k$-cycles.
The vertices in the $f^{k}$-graph that belong to $k$-cycles are called cluster heads of order k. The union of cluster heads of order k forms a sub-PP of $\Phi_{c}^{k-1}$, denoted by $\Phi^{k}_{c}$. Let $\Phi^{k}_{a}=\Phi^{0}\setminus\Phi^{k}_{c}$.
If the edges of the $k$-cycles in the $f^{k}$-graph are deleted, one obtains a collection of directed trees referred to as cluster subtrees of order k. Each vertex $x$ in a cluster subtree of order $k$ is either a cluster head of order $k$ (i.e., $x\in\Phi^{k}_{c}$) or such that $x\in\Phi^{k}_{a}$, and there is a path from $x$ to a cluster head of order k. The assumption that the intensity of $\Phi_{k-1}^{c}$ is at most $\rho/2^{k-1}$ implies that the intensity of $\Phi_{k}^{c}$ is at most $\rho/2^{k}$.
Figure 7: Left picture: the graph of $f^{3}$ with the $3$-cycles in boldface;
Right picture: Cluster subtrees of order 3.
Remark 2.6.
Denote by $\Phi_{e}^{k}$ the point process of exit points of order $k$, and let the intensity of $\Phi_{e}^{0}$ be $\rho/2$. Under the first level of the $\mathrm{CHN}^{2}$ clustering algorithm, each point in $\Phi_{e}^{0}$ is either a cluster head of order one or is connected to a cluster head of order one, and each resulting cluster contains exactly two cluster heads. Furthermore, when two cluster heads of order one belong to the same cluster of order one, one of them becomes an exit point of order one while the other does not. Therefore, the intensity of exit points of order one is less than $\rho/4$. By the same reasoning, the intensity of exit points of order $k$ is less than $\rho/2^{k+1}$. Figure 5 shows simulation estimates of the intensity of exit points at level $k$. As seen in the figure, the intensity at each level is reduced by approximately a factor of three, suggesting that the estimated intensity of exit points at level $k$ is close to $\rho/3^{k+1}$.
2.2 Construction of the Limiting Point-Shift
This subsection discusses the existence and properties of the limiting point-shift and point-shift graph of the $\mathrm{CHN}^{2}$ clustering algorithm.
Definition 2.7.
Consider the sequence $f^{n}$ of the point-shifts constructed in Subsection 2.1 on a Poisson Point Process (PPP). The fact that the intensity of $\Phi^{i}_{c}$ tends to zero as $i$ tends to infinity, as shown in Subsection 2.1, implies that the sequence $\{\Phi_{a}^{k+1}\setminus\Phi_{a}^{k}\}_{k\geq 0}$, where $\Phi_{a}^{0}=\emptyset$, forms a partition of $\Phi^{0}$. This defines a new point-shift, denoted by $f^{\infty}$. The graph of this point-shift, referred to as the $f^{\infty}$-graph and denoted by $G$, is acyclic, due to the construction of the sequence $f^{n}$.
The $f^{\infty}$ point-shift is called the Centroid Hierarchical Nearest Neighbor Point-Shift ($\mathrm{CHN}^{2}PS$), and its associated graph is called the Centroid Hierarchical Nearest Neighbor Eternal Family Forest ($\mathrm{CHN}^{2}$ EFF). When this forest is a tree, it is referred to as the Centroid Hierarchical Nearest Neighbor Eternal Family Tree ($\mathrm{CHN}^{2}$ EFT).
Corollary 2.8 (Construction of the point-shift graph).
Under the Palm probability of $\Phi^{0}$ and rooted at zero, the $\mathrm{CHN}^{2}$ EFF is the local weak limit of the $f^{n}$-graphs on $\Phi^{0}$ as $n\to\infty$. This limit graph is unimodular.
Proof.
This is assessed under the Palm probability of $\Phi^{0}$. For all balls of radius $h$ centered at 0, the restriction of $f^{n}$-graphs to this ball converges a.s. to a limit when $n$ tends to infinity. This is because the intensity of the points of $\Phi^{i}_{c}$ tends to 0 as $i$ tends to infinity. Hence, there is a random $n$ such that the restriction of the graph $G^{n}$ to this ball is fully determined by $f^{n}$ (this random integer is the largest integer $n$ such that the ball in question is empty of points of $\Phi^{n+1}_{c}$). We will denote by $[G,0]$ the limiting random graph, and by $[G^{n},0]$ the $f^{n}$-graph under the Palm probability of $\Phi^{0}$. Since for each $n$, $[G^{n},0]$ is a unimodular graph (see [5]) the limit graph $[G,0]$ is also unimodular (see [6]). ∎
Theorem 2.9.
For each $n\in\mathbb{N}$, the connected components of the graph $[G^{n},0]$ are all of type $\mathcal{F/F}$.
Proof.
In Section 5.1, it will be shown that, for all $n$, there is no infinite path in $[G^{n},0]$. This fact demonstrates that, for each $n$, all the clusters of $[G^{n},0]$ are finite. Note that the cycle indicated in the properties of the $\mathcal{F/F}$ components (see Theorem 3.10 in [6]) in each cluster of $f^{n}$-graph is the cycle of mutual nearest neighbor points, which is a cycle of length two a.s. ∎
Remark 2.10.
Theorem 2.9 implies that clusters of level $n$ in the $\mathrm{CHN}^{2}$ algorithm are almost surely finite for each $n$.
In the next proposition, it will be shown that all the connected components of the $\mathrm{CHN}^{2}$ EFF share the same distribution.
Proposition 2.11.
Let $G$ be the $\mathrm{CHN}^{2}$ EFF under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}^{d}$, and let $f^{\infty}$ be $\mathrm{CHN}^{2}PS$ on $G$. Then the distribution of $[G,0]$ is identical to the distribution of $[G,v]$, for all $v$ in the support of $\Phi^{0}$.
Proof.
It is known that there exists a directed graph $G^{\prime}$ with vertex set $\Phi^{0}$, that is constructed in a deterministic and isometry-invariant way, which is a.s. a doubly infinite path (see [17]). Consequently, a point-shift on $\Phi^{0}$ can be obtained by mapping each vertex to its subsequent vertex following the direction of this path. This point-shift is a bijection. By applying Mecke’s Point Stationary Theorem (see [5]), it can be concluded that the distribution of $[G,0]$ is identical to the distribution of $[G,v]$. ∎
Corollary 2.12.
The connected components of $\mathrm{CHN}^{2}$ EFF $[G,0]$ belong to the $\mathcal{I/I}$ class of the foil classification theorem for unimodular networks. In other words, under the Poisson Point Process (PPP), each connected component of $\mathrm{CHN}^{2}$ EFF is a one-ended tree with all its foils infinite almost surely.
Proof.
As mentioned previously, at each level $k$, the equality $\Phi^{0}=\Phi^{k}_{c}\cup\Phi^{k}_{a}$ holds, where $\Phi^{k}_{c}$ denotes the cluster heads and $\Phi^{k}_{a}$ represents the acyclic points of order $n$. For a point belonging to $\Phi^{k}_{a}$ at some level $k$, the number of its descendants is finite and does not change thereafter due to the construction of the $\mathrm{CHN}^{2}$ EFF.
Since the intensity of $\Phi^{k}_{c}$ tends to zero as $k$ approaches infinity, the sets $\{\Phi^{k+1}_{a}\setminus\Phi^{k}_{a}\}_{n\geq 0}$, with $\Phi^{0}_{a}=\emptyset$, form a partition of $\Phi^{0}$. Therefore, almost surely, there exists a level $k$ such that zero belongs to $\Phi^{k+1}_{a}\setminus\Phi^{k}_{a}$. Hence, its descendant tree is almost surely finite.
Consequently, the connected component of zero is in the $\mathcal{I/I}$ class. In other words, the cluster containing the origin in the $\mathrm{CHN}^{2}$ EFF graph is a one-ended tree.
Proposition 2.11 implies that all other connected components are also in the $\mathcal{I/I}$ class. ∎
Remark 2.13 (Evaporation and $\mathcal{I/I}$ Property).
Corollary 2.12 can also be proved using the evaporation property of the $\mathrm{CHN}^{2}$ point-shift, $f^{\infty}$, on the $\mathrm{CHN}^{2}$ EFF, $[G,0]$. Recall that a point-shift $f$ on a connected component $C\subseteq[G,0]$ evaporates $C$ if, almost surely, there is no point $x\in C$ with arbitrarily large backward iterates under $f$, i.e., $f^{-k}(x)=\emptyset$ almost surely for all sufficiently large $k$ (see [5] for more details). Corollary 24 in [5] shows that if a point-shift on a point process evaporates a connected component $C$ of its point-shift graph, then $C$ belongs to the $\mathcal{I/I}$ class. Since the intensity of exit points at level $k$ goes to zero as $k\to\infty$, it follows that $f^{\infty}$ evaporates every connected component of $[G,0]$. Hence, all such components belong to the $\mathcal{I/I}$ class.
2.2.1 The Existence of a Last Universal Common Ancestor at infinity.
Considering the genetic data of the species interpretation discussed in the introduction, one can gain clearer insights by examining the properties of the $\mathrm{CHN}^{2}$ EFF graph. Firstly, the $\mathcal{I/I}$ property of the graph shows that each species (point) belongs to an infinite phylogenetic tree. The $\mathcal{I/I}$ property also shows that there exists a LUCA (Last Universal Common Ancestor) ”at infinity”, from which all species in the cluster are descendants. This raises the question of whether there are multiple LUCAs or just one in the clustering of the Poisson point process. This is equivalent to determining the number of connected components that exist in the $\mathrm{CHN}^{2}$ EFF.
Although the number of connected components is generally unknown, in dimension $d=1$, the following theorem shows that the $\mathrm{CHN}^{2}$ EFF has only one connected component.
Theorem 2.14.
In dimension $d=1$, under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}$, the $\mathrm{CHN}^{2}$ EFF is a tree.
Proof.
To prove that $[G,0]$ contains only one connected component in dimension $1$, it suffices to show that for every fixed $r>0$, almost surely all points of $\Phi^{0}$ within the ball of radius $r$ centered at 00, denoted as $B_{r}(0)$, belong to the same component.
Recall that at each level $n$ of the $\mathrm{CHN}^{2}$ clustering algorithm, the point process $\Phi^{n}_{c}$ consists of the vertices of $f^{n}$-cycles $S_{i}^{n}$ at level $n$, (see definition in Subsection 2.1). Note that the distribution of distances between pairs of $S_{i}^{n}$ cycles, for each $i$, is stationary. Consider the following events:
$A_{n+1}=\{\text{The first $f^{n}$-cycle after point $0$ in the positive side of $\mathbb{R}$ is connected}$
$\text{to an $f^{n}$-cycle in the negative side of $\mathbb{R}$ atlevel $n+1$ of the algorithm}\}.$
Consider the two points, $x_{i}$ and $y_{i}$, which are part of the $f^{n}$-cycles belonging to $S^{n}_{i}$. Let’s define $M_{i}^{n}=\frac{x_{i}+y_{i}}{2}$, which represents the middle points of these two points. The sequence of random distances between $M_{i}^{n}$s, $\{d_{i}\}_{i\in\mathbb{Z}}=\{|M_{i}^{n}-M_{i-1}^{n}|\}_{i\in\mathbb{Z}}$, forms a stationary process. Therefore, the probability of a typical cycle connecting to its left neighbor cycle is equal to the probability of it connecting to its right neighbor cycle. Therefore,
$\sum_{n}\mathbb{P}(A_{n})=\infty.$
The Borel-Cantelli lemma implies that $A_{n}$ occurs for infinitely many $n$. Similarly, one can show that the same result holds for each n, for the event $A^{\prime}_{n}$,
$A^{\prime}_{n+1}=\{\text{The first $f^{n}$-cycle after point $0$in the negative side of $\mathbb{R}$ is connected}$
$\text{to an $f^{n}$-cycle in the positive side of $\mathbb{R}$ atlevel $n+1$ of the algorithm}\}.$
Furthermore, note that when a cluster from the right-hand side of zero connects to a cluster on the left-hand side of zero and forms a new cluster, all points between them also belong to the same cluster. It is also evident that at step $n$, the place of the first $f^{n}$-cycle from the right-hand side of zero tends to infinity as $n\to\infty$. Thus, for each $r$, there exists a sufficiently large $n$ such that the first $f^{n}$-cycle from the right-hand side of zero connects to the left-hand side, and $B_{r}(0)$ is contained in the new cluster formed by connecting these $f^{n}$-cycles to each other. ∎
The conjecture is that this graph is connected in each dimension $d$ when constructed from a PPP, i.e., there is only one LUCA in the $\mathrm{CHN}^{2}$ EFF phylogenetic tree of a PPP.
Remark 2.15.
All the results shown here for the PPP case are due to its property of not admitting a second-order descending chain (see Definition 5.1). Therefore, for any stationary point process that satisfies this property, the same results hold.
2.2.2 The Cardinality of Descendant Trees
The construction of the $\mathrm{CHN}^{2}$ EFF allows us to calculate bounds on the mean number of species within the descendant tree of a typical cluster at a specific level $k$. Note that the direction of the edges in the graph is from child to parent. The $\mathcal{I/I}$ property (of the connected component of zero), ensures a finite number of descendants for each vertex (species), so it enables us to quantify the number of species in the descendant tree.
More precisely, let the intensity of the cluster heads of level $k$, $\Phi_{c}^{k}$, be $\rho_{k}$. Then the mean of the cardinality of the descendant tree of a point belongs to $\Phi_{c}^{k}$, is $1/\rho_{k}$. It is known from Remark 2.6 that $\rho_{k}$ tend to zero, as $k$ goes to infinity.
Proposition 2.16.
Let $G$ be the $\mathrm{CHN}^{2}$ EFF under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}^{d}$. Let $N^{0}$ be the cardinality of the descendant tree of 00. Then the mean of $N^{0}$ is infinite. So $N^{0}$ is heavy-tailed.
Proof.
The rooted graph $[G,0]$ is a unimodular graph. Let $f$ be the $\mathrm{CHN}^{2}$ point-shift on $[G,0]$. Define $g[G,x,y]=\sum_{n=1}^{\infty}\mathbf{1}_{y=f^{n}(x)}$, which is well-defined and measurable. Therefore, the mass transport principle (2) holds for $g$ (see [2]).
$\mathbb{E}[\sum_{v\in V(\mathbf{G})}g[\mathbf{G},o,v]]=\mathbb{E}[\sum_{v\in V(\mathbf{G})}g[\mathbf{G},v,o]].$ (2)
The left-hand side of (2) for $g$ is equal to infinity, and its right-hand side is equal to the mean number of descendants of 00. So the mean of $N^{0}$ for $[G,0]$ is infinite. ∎
Remark 2.17.
The result stated in Proposition 2.16 remains valid for all vertices in the $\mathrm{CHN}^{2}$ EFF. This is due to Proposition 2.11, which ensures that all connected components share the same distribution.
3 Single Linkage and Hausdorff Hierarchical Nearest Neighbor Clustering Algorithm
In the $\mathrm{CHN}^{2}$ clustering algorithm defined in Sections 1 and 2, the distance between clusters is based on the centroid distance. However, alternative definitions for the distance between clusters can also be used, such as the single linkage distance which is defined for all $X,Y\subseteq\mathbb{R}^{d}$ as follows:
$d_{S}(X,Y):=\inf\left\{d(x,y):x\in X,y\in Y\right\},$ (3)
or the Hausdorff (complete linkage) distance, which is defined for all $X,Y\subseteq\mathbb{R}^{d}$ as follows:
$d_{H}(X,Y):=\max\left\{\sup_{x\in X}d(x,Y),\sup_{y\in Y}d(X,y)\right\},$ (4)
where $d(a,B):=\inf_{b\in B}d(a,b)$ quantifies the distance from a point $a$ to a subset $B$.
While using these alternative distances in the algorithm requires more processing steps than in the $\mathrm{CHN}^{2}$ algorithm, it is still possible to define each level of the algorithm in a similar way, except for the fact that only undirected graphs will be constructed here. These algorithms will only be applied to PPP.
Let $D\subset\mathbb{R}^{d}$ be a dataset. The definitions of the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) and the Hausdorff (Complete Linkage) Hierarchical Nearest Neighbor ($\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithms are as follows:
Level 0 The zero level is defined similarly to the $\mathrm{CHN}^{2}$ clustering algorithm. Each point in $D$ is connected to its nearest neighbor with an undirected edge, resulting in a collection of finite connected components (Theorem 2.1 for $n=0$) called clusters of order 0. In order to ensure that the resulting graphs do not contain cycles, only one edge of each mutual nearest neighbor pair is retained. Consequently, the resulting graph consists of clusters that form undirected trees (see Figure 8).
Figure 8: Clusters of order zero generated by the $\mathrm{SHN}^{2}$ (and $\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithm on a finite dataset. The blue edges connect each point to its nearest neighbor. A single red edge connects points that are mutual nearest neighbors. Note that the resulting graph forms an undirected forest.
Level 1 Each cluster of order 0 is connected to its nearest neighboring cluster based on the single linkage distance between clusters for the $\mathrm{SHN}^{2}$ algorithm and the Hausdorff distance for the $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithm. Specifically:
  • For the $\mathrm{SHN}^{2}$ algorithm, define the proximity point of a cluster of order 0 as the point of the cluster that is closest to its nearest neighbor cluster based on the single linkage distance (3). Connect the proximity point to its nearest neighbor in the nearest neighbor cluster.
  • For the $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithm, define the proximity point as the point within the cluster that attains the supremum in the Hausdorff distance (4) with respect to its nearest neighboring cluster. Connect the two points that attain the supremum in the Hausdorff distance (4).
This defines clusters of level 1. Since the finiteness of clusters at this level is not established here, they may be either finite or infinite. If a cluster is finite, it contains a mutual nearest neighbor pair. To ensure that the resulting graphs do not contain cycles, only one edge of each mutual nearest neighbor pair is retained. Consequently, the resulting graph consists of clusters that form undirected trees (see Figure 9).
Figure 9: Clusters of order one produced by the $\mathrm{SHN}^{2}$ clustering algorithm on a finite dataset. Each order-0 cluster is connected to its nearest neighbor cluster using the single linkage distance. The green edges indicate the new connections between clusters, while the pink points highlight the proximity points of order 0, that is, the points in each cluster that are closest to their respective nearest neighbor cluster. Note that the resulting graph forms an undirected forest.
Level $i$ ($i\geq 2$): At this stage, it is not guaranteed that the clusters of order 1 and higher are almost surely finite. Therefore, the algorithm is generalized to handle the possibility of infinite clusters, in order to extend the procedure to higher levels. At level $i$, each finite cluster of order $i-1$ is connected to its nearest neighbor cluster using either the single linkage or the Hausdorff distance, depending on the algorithm. Consequently, the notion of proximity points applies only to finite clusters at each level (see Figure 10).
In the $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithm, if the nearest neighbor of a finite cluster $C_{j}$ is the infinite cluster $C_{j}^{\prime}$ and the maximum defined in (4) is infinite, we proceed as follows: instead of directly relying on the Hausdorff distance, we identify two vertices, one in $C_{j}$ and one in $C_{j}^{\prime}$, that achieve the infimum in (3), which corresponds to the single linkage distance between the two clusters. We then connect these two vertices, and the vertex in $C_{j}$ is defined as its proximity point. By iterating this procedure, one can proceed hierarchically as above to define proximity points of order 1 or higher and clusters of order 2 or higher.
Figure 10: Clusters of order two produced by the $\mathrm{SHN}^{2}$ clustering algorithm on a finite dataset. Each finite order-1 cluster is connected to its nearest neighbor cluster using the single linkage distance. The orange edges indicate the new connections between clusters, while the cyan points highlight the proximity points of order 1, that is, the points in each order-1 cluster that are closest to their respective nearest neighbor cluster. Note that the resulting graph again forms an undirected forest.
In the following, we analyze the graphs obtained by considering these algorithms on a PPP.
Definition 3.1.
Let $\Phi^{0}$ be a Poisson point process with intensity 1, and $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (resp. $[\mathcal{G}^{n}_{\mathcal{{H}}},0]$) be the random undirected rooted graph, rooted at zero, constructed at the $n$-th level of the $\mathrm{SHN}^{2}$($\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithm under the Palm probability of $\Phi^{0}$. The vertices of $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (resp. $[\mathcal{G}^{n}_{\mathcal{{H}}},0]$) that are proximity points of level $n$ form a sub-PP of $\Phi^{0}$ denoted by $\Xi_{S}^{n}$ (resp. $\Xi_{H}^{n}$) for the $\mathrm{SHN}^{2}$ (resp. $\mathrm{H}^{2}\mathrm{N}^{2}$) clustering algorithm.
The following theorem shows that the weak limit graph of the graphs $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ exists as $n\to\infty$.
Theorem 3.2.
The weak limit graph of $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (and similarly $[\mathcal{G}^{n}_{\mathcal{{H}}},0]$) in Definition 3.1 exists as $n\to\infty$.
Proof.
See Subsection 5.2. ∎
Definition 3.3.
The weak limit graph of the random graphs $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ in Definition 3.1 is called the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) Spanning Forest and is denoted by $[\mathcal{G}_{\mathcal{{S}}},0]$. If this forest happens to be a single connected component (i.e., a tree), it is referred to as the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) Spanning Tree.
There is a close relation between the Minimal Spanning Forest (MSF) and the $\mathrm{SHN}^{2}$ Spanning Forest on a point process in $\mathbb{R}^{d}$ for $d\geq 2$. The Minimal Spanning Forest, or more precisely the Wired MSF on the point process $\Phi$, was defined by Aldous and Steele in [23] and [1] as the union of all invasion percolation trees started at every vertex in $\Phi$. Invasion percolation from $x$ grows a tree by iteratively adding the shortest available edge that connects the existing tree to a new point (see also [19]). The following corollary describes this relationship.
Corollary 3.4.
Let $\Phi^{0}$ be a Poisson point process on $\mathbb{R}^{d}$ with $d\geq 2$. The $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the MSF on $\Phi^{0}$.
Proof.
Under the Palm probability of $\Phi^{0}$, consider a fixed realization $\phi_{0}$. It suffices to show that every edge in $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ (the $n$-th step of the $\mathrm{SHN}^{2}$ clustering algorithm) also appears in the Minimal Spanning Forest (MSF) of $\phi_{0}$.
Base Case $n=0$. At step zero, the $\mathrm{SHN}^{2}$ algorithm produces the nearest neighbor graph, where each point in $\phi_{0}$ connects to its nearest neighbor. Such edges are part of any spanning tree generated by invasion percolation, and thus belong to the MSF. Hence the statement holds for $n=0$.
Inductive Step. Assume that for some $k\in\mathbb{N}$, all edges in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ are contained in the MSF of $\phi_{0}$. We need to show this also holds for $[\mathcal{G}^{k+1}_{\mathcal{{S}}},0]$.
Let $e_{xy}$ be an edge that appears for the first time at step $k+1$. By definition of the $\mathrm{SHN}^{2}$ algorithm, $e_{xy}$ connects two clusters $C_{x}$ and $C_{y}$ built at step $k$. Assume, without loss of generality, that $C_{y}$ is the nearest neighbor of $C_{x}$. By definition, $C_{x}$ is finite.
For the sake of reaching a contradiction, suppose that $e_{xy}$ is not in the MSF of $\phi_{0}$. Consider the invasion percolation tree $t_{x}^{\infty}$ rooted at $x$. This tree is infinite, but does not include the edge $e_{xy}$. A first observation is that $C_{x}$ and $t_{x}^{\infty}$ have a non-empty intersection (both contain $x$). Since $t_{x}^{\infty}$ is infinite, whereas $C_{x}$ is finite, $t_{x}^{\infty}$ contains at least one edge that connects $C_{x}$ to a point (resp. a cluster of order $k$) that does belong to $C_{x}$ (resp. different from $C_{x}$). Every such edge of $t_{x}^{\infty}$ must have length strictly less than $\lvert e_{xy}\rvert$. But this contradicts the assumption that $C_{y}$ is the nearest neighboring cluster of $C_{x}$. Thus, $e_{xy}$ must be part of the MSF.
By induction, all new edges introduced at each level $n$ also lie in the MSF. Hence $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ is a subgraph of the MSF for every $n$. Since the weak limit of $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ exists and defines the $\mathrm{SHN}^{2}$ Spanning Forest on $\Phi^{0}$, it follows that the $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the MSF on $\Phi^{0}$. ∎
Remark 3.5.
The definition of the $\mathrm{SHN}^{2}$ Spanning Forest is local, and studying its properties may provide insights into the behavior of the MSF on PPP. For example, if the $\mathrm{SHN}^{2}$ Spanning Forest is connected on a PPP in some dimension, then the MSF is also connected in that dimension.
4 Hierarchical Clustroid Nearest Neighbor Chain Algorithm on Cox Point Processes
The aim of this section is twofold. The first one is to show that the algorithms proposed here can be applied to other point processes than Poisson. The second one is to see how the algorithms in question can be used to detect data aggregates when present.
Clustering algorithms are expected to detect aggregates of points in a dataset. Working with a Poisson point process is useful in this context because the PPP can be viewed as a null hypothesis, i.e., a baseline model that has no structural underlying aggregates and where fine random local properties are leveraged to nevertheless recursively cluster data in function of the dataset only. To examine the behavior of the $\mathrm{CHN}^{2}$ algorithm in scenarios where structural aggregates do occur, a natural testbed is that of Cox point processes.
Specifically, the first example (Example 4.2) is constructed so that each realization contains distinct aggregates of points (see Remark 4.1), and aims at showing how the algorithm forms clusters when such Cox aggregates are present. The second example (Example 4.3) generalizes the first one by removing the requirement for distinct aggregates, while still only retaining a sub-point process of the PPP. As a result, the properties established in Section 2 (such as the finiteness of clusters at each hierarchical level and the weak convergence of the limiting graph) continue to hold. Moreover, it is shown in these examples how the $\mathrm{CHN}^{2}$ clustering algorithm effectively detects the presence of aggregates.
Remark 4.1.
In many clustering algorithms, the term cluster refers to a distinct subgroup of data points that are more similar to one another than to points outside that subgroup. In contrast, in this paper, the word cluster simply denotes a connected component at a given hierarchical level. To avoid confusion between these two usages, the term aggregates of points will be used in this paper to denote data with distinct groups of similar objects.
Example 4.2 (Cox point process).
A point process $N$ on a state space $\mathcal{S}$ is called a Cox point process driven by the (locally finite) random measure $M$ if, conditioned on $M$, $N$ is a Poisson point process with intensity function $M$ (see [10]).
Consider a stationary point process $\xi$, and let $r=\{r_{x}\}_{x\in\xi}$ be a real-valued random sequence of marks representing the radii associated with each point in $\xi$. Define ${\Xi}=\bigcup_{x\in\xi}B_{r_{x}}(x)$, where $B_{r_{x}}(x)$ denotes the ball of radius $r_{x}$ centered at $x$. Next, define the random measure $M$ on $\mathbb{R}^{d}$ by
$M(B)=\mu(B\cap\tilde{\xi}),$
where $\mu$ is the Lebesgue measure on $\mathbb{R}^{d}$, and $B\subseteq\mathbb{R}^{d}$ is a Borel set.
Let $N$ be the Cox point process driven by $M$. Conditioned on the realization of $\xi$, $N$ is a sub point process of a homogeneous Poisson point process. More precisely, $N$ is a Cox point process with intensity measure $M$, supported only within the random region $\tilde{\xi}$. Furthermore, due to the stationarity of $\xi$, viewed as a marked point process, $N$ is also stationary.
Since $N$ is a sub-point process of a homogeneous Poisson point process, it does not admit any second-order descending chain. Additionally, the stationarity of $N$ ensures that all results from Section 2 apply to $N$.
Figures 11 and 12 provide simulations of the $\mathrm{CHN}^{2}$ clustering algorithm on $N$, where $N$ is defined as above.
Figure 11: A realization of the Cox point process from Example 4.2, where $S$ contains three random points in $\mathbb{R}^{2}$ and $r=\{40,20,30\}$. This realization consists of approximately 2000 points, which merge into a single cluster after six steps.
Figure 12: A realization of the Cox point process from Example 4.2, where $S$ contains four random points in $\mathbb{R}^{2}$ and $r=\{30,20,30,30\}$. This realization consists of approximately 2500 points, which merge into a single cluster after seven steps.
4.1 Stopping Criteria for Detecting Data Aggregation in $\mathrm{CHN}^{2}$ Clustering
A classical question in hierarchical clustering is that of the identification of aggregates through the clustering process. It is proposed here to use stopping rules purely based on the lengths of edges used in clustering to assess the presence of aggregates. These stopping rules are defined by comparison to the null hypothesis which is that of a Poisson dataset.
To answer this, the mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm are computed and plotted for each data set in Figure 13. In this figure, the left diagram corresponds to the data set of Figure 11, while the right one corresponds to the data set of Figure 12. In both diagrams, one can observe a distinct ’jump’ in the mean distances, which can be used to determine the stopping level for clustering.
To make this more quantitative, a ratio-based jump detection method can be used. Let $d_{\text{Cox},k}$ and $d_{\text{Poisson},k}$ denote the mean nearest-neighbor distances at level $k$ for the Cox and Poisson datasets, respectively. Consider the ratio:
$R_{k}=\frac{d_{\text{Cox},k}}{d_{\text{Poisson},k}}.$
By tracking the evolution of $R_{k}$ across levels, one can identify aggregation on the basis of a sudden increases in this ratio. Specifically, aggregation is detected at the first level $k^{*}$ where the relative increase in $R_{k}$ exceeds a predefined threshold, i.e.,
$\frac{R_{k}-R_{k-1}}{R_{k-1}}>\tau.$
The results of this method are illustrated in Figure 14, where significant jumps in the mean distance ratios are observed at specific levels. For the dataset in Figure 11, the first pronounced jump occurs at level $6$, while for Figure 12, it appears at level $5$. This is for a threshold $\tau=0.3$.
Figure 13: Left Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, serving as a null hypothesis and the data set shown in Figure 11. At level 6, a pronounced jump is observed in the Cox case, suggesting visually a reasonable stopping criterion. A smaller jump is also visible at level 5, which may indicate partial detection of early-stage clusters.
Right Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, and the data set shown in Figure 12. The first significant jump is observed at level 5, suggesting visually a suitable stopping level for this aggregated dataset.
Figure 14: Left Diagram: Ratio of mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the Cox dataset shown in Figure 11. Clustering is detected at level 6 for a threshold $\tau=0.3$.
Right Diagram: Ratio of mean distances at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the dataset shown in Figure 12. Clustering is detected at level 5 for a threshold $\tau=0.3$.
The following example is a more general case of Example 4.2.
Example 4.3.
Consider $\hat{S}$, an arbitrarily defined random subset of $\mathbb{R}^{d}$. Similar to Example 4.2, define a random measure $M$ on $\mathbb{R}^{d}$ as:
$M(B)=\mu(B\cap\hat{S}),$
where $\mu$ is the Lebesgue measure on $\mathbb{R}^{d}$, and $B\subseteq\mathbb{R}^{d}$ is a Borel set.
Let $N$ be the Cox point process driven by $M$. Conditioned on the realization of $\hat{S}$, $N$ is a homogeneous Poisson point process restricted to the random set $\hat{S}$. So it does not admit any second-order descending chain. Therefore, all the results from Section 2 apply to $N$.
5 Proofs
5.1 Proof of $\mathcal{F/F}$ property of the pre-limits point-shift graphs of $\mathrm{CHN}^{2}$
In Section 2.1, we claimed that, for every $k\in\mathbb{N}$, the point-shift graph, $f^{k}$, contains no infinite path (component). This section proves this claim. For simplicity, it is first shown that the result holds for $k=1$. Then, it will be shown that the same result holds for all $k$. Before this, Second-order descending chains are introduced and it is shown in Proposition 5.3 that such chains do not exist in the Poisson point process and in Proposition 5.6 that an infinite path in the $f^{1}$-graph is a second-order descending chain.
Definition 5.1.
Let $N$ be a point process. Let $\{x_{n}\}_{n\in\mathbb{N}}$ be an infinite sequence of points in $N$ without repetition. This sequence is called a second-order descending chain if, for all $i\geq 2$, the inequality $d_{i}<\max(d_{i-1},d_{i-2})$ holds, where $d_{i}=\|x_{i+1}-x_{i}\|$.
Remark 5.2.
A descending chain $\{x_{n}\}_{n\in\mathbb{N}}$ in a point process $N$ is defined by the property that for all $i\geq 1$, the inequality $d_{i}<d_{i-1}$ holds, where $d_{i}=|x_{i+1}-x_{i}|$ (see [22] and [11]). Any descending chain in $N$ is a second-order descending chain; however, the converse does not hold.
Proposition 5.3.
A $d$-dimensional Poisson point process, $\Phi$, contains no second-order descending chain a.s.
Proof.
By classical arguments, it is enough to show that if $\Phi$ is a homogeneous PPP, the probability of having a second-order descending chain in $\Phi+\delta_{0}$, starting from 00, is zero. Consider the following events:
$A=\{\text{ there exists a second-order chain, $\{x_{n}\}_{n\in\mathbb{N}}$, in$\Phi+\delta_{0}$, with $x_{0}=0$}\},$
and
$A_{R}=\{\text{ there exists a second-order chain, $\{x_{n}\}_{n\in\mathbb{N}}$, in $\Phi+\delta_{0}$, with $x_{0}=0$, and $d_{0},d_{1}<R$}\},$
where $R$ is an integer. Since $A=\cup_{R=1}^{\infty}A_{R}$, it is enough to show that $\mathbb{P}(A_{R})=0$ for all $R$. For $n\geq 3$, let $A_{R,n}$ be the event of existence of a second-order descending chain, $\{x_{i}\}_{i=0}^{n}$ of length (number of edges) $n$, that is, a chain in $\Phi+\delta_{0}$, with $x_{0}=0$, satisfying the property $d_{i}<\max(d_{i-1},d_{i-2})$ for all $12\leq i\leq n-$ 1, and such that $d_{0},d_{1}<R$. Then one can write $A_{R}=\cap_{n\geq 3}A_{R,n}$. Since for all $i$, $A_{R,i}\supseteq A_{R,i+1}$, so
$\mathbb{P}(A_{R})=\lim_{n\to\infty}\mathbb{P}(A_{R,n}).$ (5)
Thus it is enough to show that $\lim_{n\to\infty}\mathbb{P}(A_{R,n})=0$.
For $n\geq 3$, let $X_{R,n}$ denote the number of second-order descending chains with length $n$ in $\Phi+\delta_{0}$ with $x_{0}=0$ and with $d_{0},d_{1}<R$. Since $\mathbb{P}(A_{R,n})\leq\mathbb{E}[X_{R,n}]$, it is enough to show that $\lim_{n\to\infty}\mathbb{E}[X_{R,n}]=0$.
For $n\geq 1$, consider the function $g_{n+2}(x,\Phi)$ equal to the number of second-order descending chains with length $n+2$, $x_{0}=0$, and $x_{1}=x\neq 0$, under the assumption that 0 and $x$ are in the support of $\Phi$. Then
$\mathbb{E}[X_{R,n+2}]$ $=\mathbb{E}[\,\sum_{x\in\mbox{supp}(\Phi)\cap B_{R}(0)}g_{n+2}\left(x,\Phi+\delta_{0}\right)]\,$
$=\lambda\int_{B_{R}(0)}\mathbb{E}[\,g_{n+2}\left(x,\Phi+\delta_{0}+\delta_{x}\right)]\,{\mathrm{d}}x,$
where the last equality follows from Campbell’s formula and Slivnyak’s theorem. For all $\Phi$ with a support containing $\{0\}$ and $\{x\}$, consider the function $g^{\prime}_{n+2}(y,\Phi)$ equal to the number of second-order descending chains with length $n+2$, without repeated vertices and such that $x_{0}=0$, $x_{1}=x\neq 0$, and $x_{2}=y$, with $y\neq 0$ and $y\neq x$, when assuming that $y$ is in the support of $\Phi$. By the same arguments as above,
$\mathbb{E}[X_{R,n+2}]$ $=\lambda\int_{B_{R}(0)}\mathbb{E}[\,\sum_{y\in B_{R}(x)\cap\mbox{supp}(\Phi),y\neq 0,y\neq x}g^{\prime}_{n+2}\left(y,\Phi+\delta_{0}+\delta_{x}\right)]\,{\mathrm{d}}x$
$=\lambda\int_{B_{R}(0)}\lambda\int_{B_{R}(x)}\mathbb{E}[\,g^{\prime}_{n+2}\left(y,\Phi+\delta_{0}+\delta_{x}+\delta_{y}\right)]\,{\mathrm{d}}y{\mathrm{d}}x.$
The definition of the second-order descending chain, together with the fact that we consider chains without repeated vertices and the stationarity of $\Phi$, imply that the last equation is equal to
$\lambda^{2}\int_{B_{R}(0)}\int_{B_{R}(x)}\mathbb{E}[\,X_{\max(\lVert x\rVert,\lVert y-x\rVert),n}]\,{\mathrm{d}}y{\mathrm{d}}x.$
So the following recursive equation holds for the events $X_{R,n}$:
$\mathbb{E}[X_{R,n+2}]=\lambda^{2}\int_{B_{R}(0)}\int_{B_{R}(x)}\mathbb{E}[\,X_{\max(\lVert x\rVert,\lVert y-x\rVert),n}]\,{\mathrm{d}}y{\mathrm{d}}x.$ (6)
Lemma 5.4.
Let $\Phi$ be a $d$-dimensional Poisson point process of intensity $\lambda$, and let $X_{R,n}$ be the number of second-order descending chains of length $n$ with $\max(d_{0},d_{1})<R$. Then for all $n\in\mathbb{N}$,
$\mathbb{E}[X_{R,n}]=\frac{\bigl{(}\lambda^{2}\,\omega_{d}^{2}\,R^{2d}\bigr{)}^{\lfloor n/2\rfloor}}{\lfloor n/2\rfloor!}\;\times\;\bigl{(}\lambda\,\omega_{d}\,R^{d}\bigr{)}^{\,n\bmod 2}.$ (7)
where $\omega_{d}$ is the volume of the unit $d$-dimensional ball.
Proof.
The proof is by induction on $n$. Since the recurrence relation (6) increases $n$ by $2$ at each step, there are two separate induction ”tracks”: one starting from $n=0,2,4,\dots$ and the other starting from $n=1,3,5,\dots$.
For even numbers the base case is $n=0$. The $\mathbb{E}[X_{R,0}]$, is the number of second-order descending chain with length 00 in $\Phi\cup{0}$, with $x_{0}=0$, which is equal to 1, so the base case holds for the even numbers. Assume that (7) holds for an even number $n$. It then follows from (6) that
$\mathbb{E}[X_{R,n+2}]$ $=\lambda^{2}\int_{B_{R}(0)}\int_{B_{R}(x)}\mathbb{E}[\,X_{max(\lVert x\rVert,\lVert y-x\rVert),n}]\,{\mathrm{d}}y{\mathrm{d}}x$
$=\frac{\lambda^{n+2}\omega_{d}^{n}}{n/2!}\int_{B_{R}(0)}\int_{B_{R}(x)}(\max(\lVert x\rVert,\lVert y-x\rVert))^{nd}{\mathrm{d}}y{\mathrm{d}}x.$ (8)
The inner integral is equal to
[…]
$=\frac{\omega_{d}}{n+1}(\,R^{(n+1)d}+n\lVert x\rVert^{(n+1)d}),$ (9)
where ${\mathbb{S}}^{d}$ denotes the $d$-sphere. Equation (5.1) together with (5.1) give
$\mathbb{E}[X_{R,n+2}]$ $=\frac{\lambda^{n+2}\omega_{d}^{n+1}}{(n+1)(n/2!)}\int_{B_{R}(0)}(\,R^{(n+1)d}+n\lVert x\rVert^{(n+1)d}){\mathrm{d}}x$
$=\frac{\lambda^{n+2}\omega_{d}^{n+1}}{(n+1)(n/2!)}(\,\int_{B_{R}(0)}R^{(n+1)d}{\mathrm{d}}x+\int_{B_{R}(0)}n\lVert x\rVert^{(n+1)d}{\mathrm{d}}x)\,$
$=\frac{\lambda^{n+2}\omega_{d}^{n+1}}{(n+1)(n/2!)}(\,\omega_{d}R^{(n+2)d}+\frac{n}{n+2}\omega_{d}R^{(n+2)d})\,$
$=\frac{\lambda^{n+2}\omega_{d}^{n+2}}{\frac{n+2}{2}!}R^{(n+2)d},$
which completes the proof for even numbers.
For odd $n$, the base case is $n=1$. Observe that $\mathbb{E}[X_{R,1}]$, the expected number of second-order descending chains of length $1$, is $\lambda\,\omega_{d}\,R^{d}$. This matches (7) because $\lfloor 1/2\rfloor=0$ and $(1\bmod 2)=1$. Then, by applying the same recurrence relation (6), which again advances indices in steps of two, one proves the result for all odd $n=1,3,5,\dots$. Hence the inductive argument for odd $n$ is identical to the even case, simply starting from $n=1$ instead of $n=0$. This completes the proof.
Definition 5.5.
[…]
Proposition 5.6.
Any infinite path in the $f^{1}$-graph is a second-order descending chain.
Proof.
Assume that the $f^{1}$-graph has an infinite component. Since the $f^{0}$-graph has all its components finite (Proposition 2.1), there must be an infinite path $\mathcal{P}$ on $\{S^{0}_{i}\}_{i}$ (see Definition 2.3), with a structure as depicted in Figure 15.
[…]
More precisely, let $e_{i}$ be the $i$-th edge in $\mathcal{P}$. Then $e_{i}$ can be either of $d$ or $\delta$ type. If $e_{i}$ is of the $d$-type, then, there are two possibilities for $e_{i-1}$ and $e_{i-2}$:
  1. 1.
    $e_{i-1}$ and $e_{i-2}$ are both $\delta$-type edges (see Figure 16.a). In this case, since $\delta$-type edges are consecutive edges in the descending chain on $\{S^{0}_{i}\}_{i}$, one has $\|e_{i-1}\|<\|e_{i-2}\|$. And since $e_{i}$ is a MNN0 edge, then $\|e_{i}\|<\|e_{i-1}\|$. Thus $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)$.
  2. 2.
    $e_{i-1}$ is a $\delta$-type edge and $e_{i-2}$ is a $d$-type edge (see Figure 16.b). Since $e_{i-2}$ and $e_{i}$ are both edges between MNN0 points, this gives
    $\|e_{i-1}\|>\|e_{i}\|\ \mbox{and }\|e_{i-1}\|>\|e_{i-2}\|,$ (10)
    which implies that $\|e_{i}\|<\|e_{i-1}\|=\max(\|e_{i-1}\|,\|e_{i-2}\|)$.
    Figure 16: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on $e_{i}$ is a $d^{0}$-type edge in the infinite path of the $f^{1}$-graph, if it exists.
If $e_{i}$ is a $\delta$-type edge, then there are three possibilities for $e_{i-1}$ and $e_{i-2}$.
  1. 1.
    both are $\delta$-type edges (see Figure 17.a). In this case, $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)$ since these edges are in a descending chain in $\{S^{0}_{i}\}_{i}$.
  2. 2.
    $e_{i-1}$ is a $d$-type edge and $e_{i-2}$ is a $\delta$-type edge (see Figure 17.b). Then
    $\|e_{i-2}\|>\|e_{i}\|\quad\mbox{and}\quad\|e_{i-2}\|>\|e_{i-1}\|,$ (11)
    where the first inequality follows from the descending chain property and the second one holds since $e_{i-1}$ is a MNN0 edge between points. This implies that $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)=e_{i-2}$.
  3. 3.
    $e_{i-1}$ is a $\delta$-type edge and $e_{i-2}$ is a $d$-type edge (see Figure 17.c). It is then easy to show that, in this case also, the inequality $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)$ holds.
This completes the proof of Proposition 5.6.
Figure 17: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on $e_{i}$ is a $d$-type edge in the infinite path of the $f^{1}$-graph if it exists.
Propositions 5.3 and 5.6 give
Corollary 5.7.
If $\Phi^{0}$ is a Poisson point process, then the components of $f^{1}$-graph of $\Phi^{0}$ all a.s finite.
So far, it has been shown that there is no infinite path or cluster in the $f^{1}$-graph. The following proposition establishes that this holds true for all $f^{n}$-graphs.
Proposition 5.8.
For all $n\in\mathbb{N}$, any infinite path in $f^{n}$-graph is a second-order descending chain.
Proof.
The proof is by induction. Any infinite path in the $f^{0}$-graph is a descending chain in the PPP. Therefore, the base case of the induction is true. For the induction step, suppose there is no infinite path (cluster) in the $f^{n}$-graph. Then, if an infinite path $P^{n+1}$ exists in the $f^{n+1}$-graph, this path lies on $\{S^{n+1}_{i}\}_{i}$ eventually. Thus, there are two types of edges on $P^{n}$: the first type is that between two ${\mathrm{MNNn}}$ points, and the second type is that of edges that connect an exit point of level $n$, $x$, to ${\mathrm{NN(n+1)}}(x)$. The construction of the $f^{n+1}$ point-shift shows that $P^{n+1}$ forms a descending chain between the $f^{n}$-cycles $\{S_{n}^{i}\}$. The same proof as in Proposition 5.6 shows that if $P^{n}$ exists, it is a second-order descending chain. ∎
5.2 Proof of Theorem 3.2
To prove Theorem 3.2, an approach similar to that of the proof of Corollary 2.8 is proposed. We recall that the proof of Corollary 2.8 relies on the fact that the intensity of the exit points at level $k$ of the $\mathrm{CHN}^{2}$ clustering algorithm, tends to zero as $k\to\infty$. For the weak convergence of the graphs $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ constructed at level $k$ in the $\mathrm{SHN}^{2}$ clustering algorithm, a similar argument is required. Specifically, it must be shown that the intensity of proximity points of level $k$, denoted by $\lambda_{k}$, also tends to zero as $k\to\infty$. To establish this, this section defines a sequence of vertex shifts, $f^{k}_{s}$, for $k\geq 0$, on the random graphs $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$. The definition of these vertex shifts is based on the notion of mutual nearest neighbor pairs, which are well defined within each finite cluster. Head points and exit points are defined as in Section 2 for the $\mathrm{CHN}^{2}$ algorithm. Finally, using the intensity of exit points, it is shown that the intensity of proximity points at each level tends to zero as $k\to\infty$.
It is important to note that the vertex shift of order $k$ does not coincide with the edges of the graph at different levels of the algorithm. In particular, the vertex shift graph on $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ is only used for assessing $\lambda_{k}$, and it should not be confused with the undirected graph of interest here.
Order 0
[…]
Each connected component of the $f^{0}_{s}$-graph on the PPP contains a cycle of length two. These cycles are again referred to as 0-cycles. The vertices in the $f^{0}_{s}$-graph that belong to the 00-cycles are still called cluster heads of order zero. The union of cluster heads of order zero forms a thinning of $\Phi^{0}$, denoted by $\Phi^{0}_{\mathcal{S},c}$. Let $\Phi^{s,0}_{a}=\Phi^{0}\setminus\Phi^{s,0}_{c}$, which denotes points not in 00-cycles.
Order 1
To define the second point-shift, mutual nearest neighbors within the clusters of order zero are used. The definition is similar to the $\mathrm{CHN}^{2}$ clustering algorithm on PPP, but instead of using the distance between mutual nearest neighbor pairs within clusters, we use the distance between clusters, specifically with respect to the single linkage distance. For this, we need the following definition:
Definition 5.9.
Let $\{C^{0}_{i}\}_{i}$ denote the collection of 00-cycles, and let $\{S^{0}_{i}\}_{i}$ be the collection of pairs of points in these cycles. Specifically, $S^{0}_{i}$ is the pair of head points belonging to the cycle $C^{0}_{i}$, and $\mathcal{C}^{0}_{i}$ is the cluster in $[\mathcal{G}^{1}_{\mathcal{S}},0]$ that $C^{0}_{i}$ belongs to. The sets $\{S^{0}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{s,0}_{c}$.
Now, consider the nearest neighbor map, denoted by NN1, on $\{S^{0}_{i}\}_{i}$. This map associates each pair $S^{0}_{i}$ with $S^{0}_{j}$, where $S^{0}_{j}$ belongs to $\mathcal{C}^{0}_{j}$, and $\mathcal{C}^{0}_{j}$ is the nearest neighbor of $\mathcal{C}^{0}_{i}$ based on the single linkage distance.
[…]
Define $f^{1}_{s}$ to be the vertex shift on the vertices of $[G^{1}_{S},0]$, where $f^{1}_{s}$ coincides with $f^{0}_{s}$ everywhere except at the exit points of order 0. For each such point, say $x\in S^{0}$, where $S^{0}$ is the pair in which $x$ is the exit point, define $f^{1}_{s}(x)=\mathrm{nn1}(x)$.
[…]
Definition 5.10.
[…]
If the edges of the $1$-cycles in the $f^{1}_{s}$-graph are removed, the remaining structure consists of directed trees, which we call cluster subtrees of order 1. Each vertex $x$ in a cluster subtree of order 1 is either a cluster head of order 1 (i.e., $x\in\Phi^{s,1}_{c}$) or an element of $\Phi^{s,1}_{a}$, with a directed path leading from $x$ to a cluster head of order 1. This structure is analogous to the point-shift graph of order 1 in the $\mathrm{CHN}^{2}$ clustering algorithm on Poisson point processes.
Order 2
To define the next vertex shift, $f^{1}_{s}$, on the graph $[\mathcal{G}^{2}_{\mathcal{S}},0]$, we need to introduce some concepts similar to those in Definition 5.9.
Definition 5.11.
Let $\{C^{1}_{i}\}_{i}$ denote the collection of $1$-cycles, and let $\{S^{1}_{i}\}_{i}$ represent the collection of point pairs in these cycles. Specifically, $S^{1}_{i}$ is the pair of head points in the cycle $C^{1}_{i}$, and $\mathcal{C}^{1}_{i}$ is the connected component within $[\mathcal{G}^{2}_{\mathcal{S}},0]$ to which $C^{1}_{i}$ belongs. The sets $\{S^{1}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{s,1}_{c}$.
Now, consider the nearest neighbor map, denoted by NN2, on $\{S^{1}_{i}\}_{i}$. This map associates each pair $S^{1}_{i}$ with $S^{1}_{j}$, where $S^{1}_{j}$ belongs to $\mathcal{C}^{1}_{j}$, and $\mathcal{C}^{1}_{j}$ is the nearest neighbor of $\mathcal{C}^{1}_{i}$ based on the single linkage distance.
[…]
Define $f^{1}_{s}$ as the vertex shift on $[\mathcal{G}^{2}_{\mathcal{S}},0]$. Note that $[\mathcal{G}^{2}_{\mathcal{S}},0]$ coincides with $[\mathcal{G}^{1}_{\mathcal{S}},0]$ except for the proximity points of each connected component of $[\mathcal{G}^{1}_{\mathcal{S}},0]$. Define $f^{1}_{s}$ coincides with $f^{1}_{s}$ everywhere except for the exit points of order 1. For each such point, say $x\in S^{2}$, where $S^{2}$ is the pair in which $x$ is the exit point, define $f^{2}(x)=\mathrm{nn2}(x)$. Again, $f^{2}$ defines a vertex shift on $[\mathcal{G}^{2}_{\mathcal{S}},0]$. The connected components of the $f^{2}$-graph are identical to those of $[\mathcal{G}^{2}_{\mathcal{S}},0]$, though the edges do not coincide everywhere. As mentioned earlier, without loss of generality, we assume that the connected components of the $f^{1}_{s}$-graph may be finite or infinite with positive probability. The finite connected components contain a unique cycle of length 2 almost surely, as ties do not occur almost surely. The points within these 2-cycles are exit points of order 1 and form a sub-Poisson point process (sub-PP) of $\Phi^{s,1}_{c}$, called the cluster heads of order 2 and denoted by $\Phi_{c}^{S,2}$. Let $\Phi_{a}^{S,2}:=\Phi^{0}\setminus\Phi_{c}^{S,2}$, representing the vertices in the acyclic part of the $f^{2}$-graph.
Order k
Similarly, we can define, by induction, analogous notions for each level $k\geq 2$:
As illustrated in the definitions of point-shifts $f^{1}_{s}$ and $f^{1}_{s}$, the point-shift $f^{k}_{s}$ at level $k$ can be constructed using the previous level’s point-shift, $f^{k-1}_{s}$. By induction, for each level $k$, we define the following notions:
$k$-cycles
$\{C^{k}_{i}\}_{i}$, consisting of connected components of $f^{k}_{s}$ graphs that contain cycles of length 2;
Pairs within $k$-cycles
$\{S^{k}_{i}\}_{i}$, where each pair in $\{S^{k}_{i}\}_{i}$ represents the head points of the cycle $C^{k}_{i}$ in the $f^{k}_{s}$-graph;
Nearest Neighbor Map, NNk
: for each pair $S^{k}_{i}$, $\mathrm{NNk}(S^{k}_{i})$ denotes the nearest neighboring pair $S^{k}_{j}$ in the single linkage sense, where $S^{k}_{j}$ belongs to the nearest neighbor cluster $\mathcal{C}^{k}_{j}$ of $\mathcal{C}^{k}_{i}$;
Nearest Neighbor Point Map, $\mathbf{nnk}$
: for each pair $S^{k}_{i}$, $\mathrm{nnk}(x_{i})=x_{j}$ identifies the nearest neighboring point $x_{j}\in S^{k}_{j}$ for a point $x_{i}\in S^{k}_{i}$, achieving the minimum distance between the two pairs, when $\mathrm{NNk}(S^{k}_{i})=S^{k}_{j}$.
Points within these $k$-cycles which are exit points of order $k-1$ form a sub point process (sub-PP) of $\Phi^{s,k}_{c}$, referred to as cluster heads of order $k$ and denoted by $\Phi_{c}^{s,k}$. Let $\Phi_{a}^{s,k}:=\Phi^{0}\setminus\Phi_{c}^{s,k}$ represent the vertices in the acyclic part of the $f^{k}_{s}$-graph.
Remark 5.12.
Let $\Phi^{s,k}_{e}$ be the point process of exit points of order $k$, obtained by applying the vertex shift $f^{k}_{s}$. Suppose the intensity of $\Phi^{s,0}_{e}$ is $\rho$. By the same argument given in Remark 2.6, one can show that the intensity of $\Phi^{s,k}_{e}$ is less than $\rho/2^{k}+1$. Consequently, the weak limit graph of the vertex-shift graphs ($f^{k}_{s}$-graphs) exists and is denoted by the $f^{\infty}_{s}$ graph.
Proof of Theorem 3.2.
Consider the $\mathrm{SHN}^{2}$ (or $\mathrm{H}^{2}\mathrm{N}^{2}$) algorithm on $\Phi^{0}$. By the definition of $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ and of $f^{k}_{s}$-graphs, each connected component of $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ contains exactly the same vertices as the corresponding connected component of the $f^{k}_{s}$-graph, and vice versa. That is, each connected component $C_{k}$ in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ can be mapped in a bijective way to one connected component $C^{\prime}_{k}$ in the $f^{k}_{s}$-graph.
Similarly, each proximity point in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ can be mapped bijectively to one exit point in the $f^{k}_{s}$-graph, and vice versa. Hence, the intensity $\lambda_{k}$ of the proximity points in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ equals the intensity of the order-$k$ exit points in the $f^{k}_{s}$-graph. By Remark 5.12, one concludes that $\lambda_{k}<\rho/2^{k}$.
Since $\lambda_{k}\to 0$ as $k\to\infty$, the same argument as the one used in the proof of Corollary 2.8 applies here to show that the weak limit of the graphs $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ on $\Phi^{0}$ exists as $n\to\infty$. ∎
6 Connection with Other Spanning Forests on Poisson Point Process
In this paper a hierarchical clustering model is introduced using the $\mathrm{CHN}^{2}$ clustering algorithm, which creates a sequence of random graphs in a Poisson point process (PPP). Under Palm probability, its weak limit graph, $\mathrm{CHN}^{2}$ Spanning Forest, exists. The $\mathrm{CHN}^{2}$ Spanning Forest is a forest on PPP, where each connected component of it is one-ended. This section aims to compare the similarities and differences between this model and other known structures in the literature constructed on PPP.
The first construction related to $\mathrm{CHN}^{2}$ Spanning Forest is the factor graphs on PPP. Ferrari, Landim, and Thorisson [14] provide a deterministic algorithm to construct a graph with no loops (either a tree or a forest), where the vertices are the points of a d-dimensional stationary Poisson process. For dimensions 2 and 3, their algorithm results in a one-ended tree. Holroyed and Peres [17] shown that in any dimension $d$, a factor graph of the Poisson point process $\Phi^{0}$ can be constructed as a one-ended tree. A factor graph on a point process is a random graph $G$ whose vertex set is the support of the point process, and $G$ is a deterministic function invariant under isomorphisms on the PPP. In this sense, the $\mathrm{CHN}^{2}$ EFF graph $[G,0]$ defined on the PPP can also be seen as a factor graph, forming a one-ended forest.
In their proof, Holroyed and Peres construct different levels $n\in\mathbb{N}$ of coarser partitions on the PPP, resulting in infinitely many finite connected components at each level. Each component at level $n$ belongs to exactly one component at level $n+1$, and in the limit, there remains only one connected component. Using this structure, they define a one-ended tree on the PPP. While their construction involves a hierarchical structure similar to ours, it is analytical and does not provide an explicit algorithm for constructing the different levels of connected components.
As an extension, Timar [25] showed that the existence of a tree factor graph on the PPP can be extended to any stationary non-equidistant point process.
Another well-known model is the Minimal Spanning Forest (MSF). Aldous and Steele [23] introduced the MSF as a natural extension of the Minimal Spanning Tree (MST) to infinite graphs and point processes. They conjectured that for the PPP, their MSF consists of a single one-ended tree. However, it remains generally unknown how many components this algorithm produces and how many ends each component has (some results for dimension 2 are found in Alexander’s work [3]). Notably, the first level of the $\mathrm{CHN}^{2}$ clustering algorithm, i.e., $[G^{0},0]$, forms a subgraph of the MSF on the PPP. However, from the second level onward, due to the algorithm’s greedy nature, significant differences emerge.
Finally, a model related to random thinning, which naturally leads to a classification algorithm, is another relevant structure. Consider a stationary point process $\Phi$. Let $\pi=\{p_{k}\}$ be a probability distribution on the positive integers. Independently mark the points of $\Phi$ according to $\pi$. Then, perform a random thinning of $\Phi$, retaining only points of $\Phi^{0}$ marked by $k$. Declare the retained points $\Phi^{k}_{c}$ to be the cluster heads of level $k$, and assign as descendants the points of $\Phi^{k-1}_{c}$ that lie within the Voronoi cell of their nearest cluster head in $\Phi^{k}_{c}$. This algorithm, studied in [27], requires additional randomness, while clustering in $\mathrm{CHN}^{2}$ does not.
7 Open Problems
Several open questions and conjectures remain regarding the $\mathrm{CHN}^{2}$ point-shift and the $\mathrm{CHN}^{2}$ EFF. One such question revolves around whether the $\mathrm{CHN}^{2}$ EFF is a tree when constructed on the PPP. The conjecture is that, in all dimensions, this point-shift graph is a tree on a PPP.
Consider a point-shift $f$ on a point process (PP) and the action of the semigroup of translations by $-f$ on probability distributions of counting measures that have a point at the origin. The $f$-probability (point map probability) of the point process is then defined as the limit of the orbit of this semigroup action on the Palm distribution of the PP when it exists. Another question related to $\mathrm{CHN}^{2}$ point-shift is whether the point map probability of this point-shift exists on the PPP, and how to compute it if it exists. Answering these questions and validating the conjecture would greatly contribute to our understanding of the algorithm’s behavior.
In the context of the $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ clustering algorithms on Poisson point processes (PPP), a number of questions remain open as well:
  • In Corollary 3.4, it is shown that on a PPP the $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the Wired MSF. While these two graphs coincide for any finite point set, an open question is whether they are also identical on a PPP.
  • Are all clusters at different levels of the algorithm almost surely finite in $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ clustering algorithms?
  • How many ends does the $\mathrm{SHN}^{2}$ Spanning Forest have in theses cases?
  • How many connected components does the $\mathrm{SHN}^{2}$ Spanning Forest have in these cases? In particular, is it a single tree or a forest?
Acknowledgements
This work was supported by the ERC NEMO grant, under the European Union’s Horizon 2020 research and innovation programme, grant agreement number 788851 to INRIA.