### v3

# Hierarchical Clustering Algorithms on Poisson and Cox Point Processes

Abstract
This paper introduces a hierarchical clustering algorithm, the Clustroid Hierarchical Nearest Neighbor ($\mathrm{CHN}^{2}$), designed for datasets with a countably infinite number of points. The method builds clusters across successive levels by linking nearest-neighbor points or clusters using the clustroid distance. The properties of this algorithm make it suitable for very large datasets.
To evaluate its properties, we first apply the algorithm to the homogeneous Poisson point process, which serves as a natural null-hypothesis model with no intrinsic aggregation. In this setting, the algorithm generates a random forest that is a factor of the Poisson point process and hence unimodular. We prove that at every level, the level-$k$ graph has only finite connected components (a.s.) and derive bounds on their mean size. We also establish the existence of a limiting graph as the number of levels tends to infinity. In this limit, clusters are infinite and one-ended, which induces a natural order within each component and supports a tree-like phylogenetic interpretation.
Beyond the Poisson case, we extend the analysis to a class of Cox and more general stationary point processes without second-order descending chains (introduced here), for which analogous results hold. Simulations show that comparing these cases with the Poisson baseline allows an efficient detection of aggregation, thereby linking the stochastic-geometric analysis to practical clustering tasks.
keywords:
Hierarchical Clustering , Clustering on Poisson Point Process , Cox Point Process , Random Trees , Unimodular Random Graphs
[…]
[1]organization=INRIA/ENS Paris
1 Introduction
Clustering is a core task in unsupervised learning: given a representation of objects, one seeks a structure formed by a similarity notion. Classical methods often target moderate-sized finite datasets. In contrast, this work introduces and analyzes a hierarchical distance–based algorithm that applies equally well to very large data sets and even to countably infinite point sets.
We begin with a brief reminder of hierarchical clustering. One starts from singleton components, specifies a notion of proximity between components, and repeatedly links the closest ones to form larger components. The outcome is a multiscale summary of the data. Common choices for between-component proximity include minimum interpoint distance (single linkage), maximum/Hausdorff distance (complete linkage), and centroid- or clustroid-based variants; see [Murtagh, Murtagh2]. The key modeling choice is how components are represented when they interact.
Our focus is on the following algorithm: the Clustroid Hierarchical Nearest Neighbor algorithm, denoted by $\mathrm{CHN}^{2}$. The name reflects two design decisions. First, each component is represented by a clustroid, i.e., a representative chosen from the data rather than an external centroid. Second, components are linked using a nearest-neighbor rule (the superscript “22” is shorthand for “NN,” so we write $\mathrm{CHN}^{2}$ instead of $\mathrm{CHNN}$). Informally, the algorithm works as follows. At level 0, every point connects to its nearest neighbor. This produces finite components organized around mutual nearest-neighbor (MNN) pairs; the two points in such a pair form the component’s clustroid pair. Under mild general-position assumptions (no distance ties) this pair is a.s. unique. At higher levels, each current component is summarized solely by its clustroid pair, and components are linked to their nearest other component with respect to the clustroid distance (the distance between clustroid pairs). The connected components of the resulting graph are the clusters at that level. Iterating this rule builds a hierarchy. Figure 1 illustrates the level-0 picture on a finite dataset; formal definitions are deferred to Section 2.
Figure 1: Clusters of order zero generated by the $\mathrm{CHN}^{2}$ clustering algorithm on a finite dataset. The blue arrows connect each point to its nearest neighbor. The red cycles represent cycles of order zero, connecting points that are mutual nearest neighbors. The details of the elements within one cluster are shown in the figure.
A simple phylogeny picture is useful here. One can view clustroid pairs as archetypes of their components. Linking archetypes across levels organizes the data into a tree-like summary: fine distinctions appear near level 0, while broader relations appear at higher levels. This finite-data interpretation guides the intuition; a rigorous asymptotic counterpart will appear later when we study the Poisson model.
The design of the algorithm scales well to large data. Once a component forms, only its two clustroid points are needed to decide future links. There is no extra optimization step to find representatives—the clustroid pair arises automatically as an MNN pair inside the component. This autonomy reduces computation and memory, since decisions are made on representatives rather than on all points, and it also enables parallelization: widely separated regions can be processed independently and hence in parallel.
We now fix the terminology. In what follows, clusters are the connected components produced by the algorithm at a given level, whereas aggregation refers to distinguishable group structure in the data itself (in the sense used in data analysis).
To study properties at the scale of very large or infinite datasets, we place the algorithm in a point-process framework. As a baseline, we adopt a null model with no intrinsic aggregation and apply $\mathrm{CHN}^{2}$ to the homogeneous Poisson point process (PPP). It will be shown that, in this setting, the algorithm produces a random forest that is a factor of the Poisson point process and hence unimodular (so the mass-transport arguments apply; see [Aldous] for more details). Unimodularity is important because it turns global properties into local expectations and gives structural leverage: for instance, in a unimodular random graph, the Foil Classification Theorem (see page 10 in [Baccelli2017]) classifies the size and the number of ends in the clusters. In our case, we prove that, at every level, all clusters are almost surely finite; we derive bounds on their mean sizes. We also establish the existence of a limiting graph as the number of levels tends to infinity. we prove that the mean size of the typical cluster in this limiting graph is infinite and all the clusters are one-ended. Returning to the phylogeny picture, one-endedness yields a natural order within the data (species) in each infinite cluster.
It is worth emphasizing that these finiteness and structural results are far from trivial in the infinite setting. In finite datasets, clusters produced by hierarchical algorithms are necessarily finite, but when the data form a stationary point process, the situation is subtler. Even the nearest-neighbor graph (the first level of the hierarchy) may contain infinite components; this occurs, for example, in certain non-Poisson stationary processes that admit infinite descending chains (see [Daley]). The proof that the homogeneous Poisson process yields only finite clusters thus requires specific geometric arguments and plays a key role in establishing the well-definedness of the hierarchy at all levels. Likewise, in the limiting infinite graph, the fact that all connected components are one-ended is not automatic: graphs with multiple ends can exhibit branching structures where no consistent order can be defined between distant points. The one-endedness property therefore ensures that, even in the infinite-data regime, each cluster admits a natural hierarchical order—providing the basis for its phylogenetic interpretation.
The main new methodological tool is the notion second-order descending chain introduced here. We show that the PPP admits almost surely no infinite second-order descending chain. This is instrumental in proving the finiteness of the level-$k$ clusters and yields intensity bounds for the level-$k$ representatives.
The approach is extended beyond the Poisson case. Cox point processes serve as prototypical models with aggregation. We introduce a class of Cox point processes for which all results established in the Poisson case hold. More generally, the results extend to all stationary point processes without second-order descending chains. In these settings, $\mathrm{CHN}^{2}$ has the same finiteness properties as the Poisson case. Simulations on the Cox processes show that, relative to the Poisson baseline, the algorithm efficiently detects aggregation, thereby distinguishing structural organization from purely stochastic dispersion.
Our contributions are the following: (i) the definition and study $\mathrm{CHN}^{2}$, a hierarchical nearest-neighbor algorithm that compresses components to clustroid pairs and scales to very large and countably infinite datasets; (ii) on the PPP, the proof of almost-sure finiteness at every level, bounds on mean cluster sizes, unimodularity of the level graphs, and the existence of a unimodular one-ended limiting graph; (iii) the introduction of the second-order descending chain method (iv) intensity bounds for representatives; and (v) the extention of the analysis to a class Cox and other stationary point processes without second-order descending chains (vi) the evidence for aggregation detection via simulations relative to the Poisson baseline.
The paper is organized as follows. Section 2 introduces the $\mathrm{CHN}^{2}$ algorithm on the homogeneous Poisson point process, defining its level-by-level construction through a sequence of point-shifts and the corresponding hierarchical clusters. This section also establishes the scale invariance of the construction, the finiteness property of each pre-limit graph, and the existence and unimodularity of the limiting graph, referred to as the $\mathrm{CHN}^{2}$ Eternal Family Forest (EFF). Section 3 extends the $\mathrm{CHN}^{2}$ clustering framework to Cox and other stationary point processes, illustrating its use for detecting point aggregation through numerical simulations and proposing stopping criteria for identifying hierarchical levels of aggregation. Section 4 provides the proofs of finiteness for all pre-limit point-shift graphs, showing that they contain no infinite paths by excluding the existence of second-order descending chains. Finally, Section 5 discusses several open questions and conjectures related to the connectivity and ergodic properties of the limiting graph, as well as the possible existence of a point-map probability for the $\mathrm{CHN}^{2}$ point-shift.
2 Clustroid Hierarchical Nearest Neighbor Clustering on a Point Process
This section considers the $\mathrm{CHN}^{2}$ algorithm on the homogeneous PPP. As previously stated in the introduction, each level of the algorithm can be characterized as a point-shift on the PPP. A point-shift $f$ maps, in a translation invariant way, each point of a stationary point process $\phi$ to some point of $\phi$ (see [Mecke], [Tho00], and [Pointshift] for more details).
The homogeneous PPP in $\mathbb{R}^{d}$ is denoted by
$\Phi^{0}=\sum_{i}\delta_{x_{i}}.$
The construction described below is scale-invariant, so there is no restriction in assuming that its intensity is $1$.
2.1 Construction of the Pre-limit Point-Shifts
We construct a sequence $f^{n}$, $n\geq 0$, of point-shifts on $\Phi^{0}$ by induction. For this, we will use the following pseudo-distance function acting on two sets of points in $\mathbb{R}^{d}$:
$\delta(S,T)=\min_{x\in S,\,y\in T}\mathfrak{d}(x,y),$ (1)
where $\mathfrak{d}$ denotes the Euclidean distance. As mentioned in the introduction, this pseudo-distance is called the single linkage distance in the literature. This function is non-negative and symmetric, but it does not satisfy the triangle inequality. Additionally, different sets have a pseudo-distance of zero whenever they share at least one point.
Preliminaries on unimodularity.
Before proceeding, let us recall that a unimodular graph is a random rooted graph $[G,o]$ that satisfies the Mass Transport Principle (MTP):
$\mathbb{E}\Big[\sum_{x\in V(G)}f(G,o,x)\Big]=\mathbb{E}\Big[\sum_{x\in V(G)}f(G,x,o)\Big],$
for all nonnegative measurable functions $f$ defined on doubly rooted graphs (see [Aldous, Baccelli2017]). Unimodularity ensures that global statements can be derived from local information. As mentioned in the introduction, this property is repeatedly used in our proofs. It allows us to apply general results such as the foil classification theorem for unimodular networks, and the NoInfinite/Finite Inclusion Theorem (see [Baccelli2017, Aldous] for details).
Foils. Given a point-shift $f$ on a (random) rooted graph $[G,o]$, define the equivalence relation $x\sim y\iff\exists\,m,n\geq 0\text{ such that }f^{m}(x)=f^{n}(y)$. The equivalence classes for $\sim$ are called foils (they can be viewed as level sets transverse to the flow induced by the parent map $f$). The Foil Classification Theorem [Baccelli2017] classifies components according to the finiteness/infiniteness of the component and of its foils (e.g., $\mathcal{F/F}$, $\mathcal{I/I}$).
Order 0
The first point-shift, $f^{0}$, is the nearest-neighbor point-shift on $\Phi^{0}$ with respect to the Euclidean distance $\mathfrak{d}$ between points. That is, for all $x\in\Phi^{0}$, we define $f^{0}(x)=\mathrm{NN0}(x)$, where $\mathrm{NN0}$ maps a point to its nearest neighbor in $\Phi^{0}$.
Proposition 2.1.
The graph of the point-shift $f^{0}$ (the $f^{0}$-graph) is unimodular, and all its connected components belong to the $\mathcal{F/F}$ class in the sense of [Baccelli2017]; that is, each component is almost surely finite and contains exactly one cycle. Equivalently, both the component and its foils (equivalence classes induced by the point-shift) are a.s. finite.
Proof.
The unimodularity of the $f^{0}$-graph follows from the fact that it is obtained as a covariant shift on a stationary point process under the Palm probability (see [Baccelli2017, Aldous]). To see that all the connected components of this graph is finite, note that any infinite path in the $f^{0}$-graph on the PPP would form a descending chain, i.e., an infinite sequence of points whose consecutive distances form a decreasing sequence (see [Meester, Matching, Daley]). However, it is shown in [Daley] that the homogeneous PPP does not admit any descending chain almost surely. Therefore, the connected components of the $f^{0}$-graph are almost surely finite and consist of directed trees attached to a single cycle connecting two mutual nearest neighbors of $\Phi^{0}$. ∎
Definition 2.2.
Each connected component of the $f^{0}$-graph on the PPP has a cycle of length two. These cycles are called 0-cycles. The connected components of the $f^{0}$-graph are also referred to as clusters of order zero. The vertices in the $f^{0}$-graph that belong to 0-cycles are called cluster heads of order zero. The union of cluster heads of order zero forms a sub-PP of $\Phi^{0}$ denoted by $\Phi^{0}_{c}$. Let $\Phi^{0}_{a}=\Phi^{0}\setminus\Phi^{0}_{c}$.
If the edges of the 0-cycles in the $f^{0}$-graph are deleted, one obtains a collection of directed trees referred to as the cluster subtrees of order 0. Each vertex $x$ in a cluster subtree of order 0 is either a cluster head of order zero (i.e., $x\in\Phi^{0}_{c}$) or is such that $x\in\Phi^{0}_{a}$, and there is a path from $x$ to a cluster head of order 0. See Figure 2.
Figure 2: Left picture: The $f^{0}$-graph generated on PPP in blue with the 0-cycles shown in bold;
Right picture: Cluster subtrees of order 0 obtained by deleting the edges of the 0-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order zero, and the other vertices have directed paths towards the cluster head of order zero within the cluster.
Order 1
Before proceeding to the definition of the point-shift $f^{1}$, we need the following definitions:
Definition 2.3.
Let $\{C^{0}_{i}\}_{i}$ denote the collection of 0-cycles, and let $\{S^{0}_{i}\}_{i}$ be the collection of pairs of points in these cycles, namely $S^{0}_{i}$ is the pair of head points belonging to the cycle $C^{0}_{i}$. The sets $\{S^{0}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{0}_{c}$.
Consider the nearest neighbor map, denoted by NN1, on $\{S^{0}_{i}\}_{i}$, which maps each pair $S^{0}_{i}$ to its nearest neighbor pair in $\{S^{0}_{j}\}_{j}$. The distance between the sets is considered to be the single linkage distance, $\delta$, defined in (1). Specifically, $S_{i}^{0}$ maps to $S_{j}^{0}$ if $\delta(S_{j}^{0},S_{i}^{0})<\delta(S_{k}^{0},S_{i}^{0})$ for all $k\neq i,j$.
The NN1 graph is a directed graph on $\{S^{0}_{i}\}_{i}$. There is an edge from $S^{0}_{i}$ to $\mathrm{NN1}(S^{0}_{i})$, the nearest neighbor of $S^{0}_{i}$ based on the single linkage distance. Let $\mathrm{NN1}(S^{0}_{i})=S^{0}_{j}$. There is a point, denoted $x_{i}$, in $S^{0}_{i}$ and a point, $x_{j}$, in $S^{0}_{j}$ that achieve the minimum in $\delta(S^{0}_{i},S^{0}_{j})$, denoted by $nn1(x_{i})=x_{j}$. The point $x_{i}$ is called the exit point of order zero in the pair $S^{0}_{i}$.
Define $f^{1}$ to be the point-shift on $\Phi^{0}$ that coincides with $f^{0}$ everywhere except for the exit points of order 0. For each such point, say $x\in S^{0}$, where $S^{0}$ is the pair where $x$ is the exit point, define $f^{1}(x)=\mathrm{nn1}(x)$. In other words, for each exit point $x$ of an $f^{0}$-cycle, $f^{1}$ is obtained from $f^{0}$ by replacing the initial image (which was the mutual nearest neighbor of $x$ with respect to $d$) by $\mathrm{nn1}(x)$.
Definition 2.4.
Each connected component of the graph of $f^{1}$ on $\Phi^{0}$ is referred to as a cluster of order 1. It will be shown in Corollary 4.6 that these connected components are almost surely $\mathcal{F/F}$, meaning that each cluster in the graph of $f^{1}$ is finite and contains a unique cycle. These cycles are of length 2 a.s. and connect two exit points of order 0, denoted $x$ and $y$, such that $\mathrm{nn1}(x)=y$ and $\mathrm{nn1}(y)=x$. These cycles are called $1$-cycles.
The vertices in the $f^{1}$-graph that belong to $1$-cycles are called cluster heads of order 1. The union of cluster heads of order 1 forms a sub-PP of $\Phi_{c}^{0}$, denoted by $\Phi^{1}_{c}$. Let $\Phi^{1}_{a}=\Phi^{0}\setminus\Phi^{1}_{c}$.
If the edges of the $1$-cycles in the $f^{1}$-graph are deleted, one obtains a collection of directed trees referred to as cluster subtrees of order 1. Each vertex $x$ in a cluster subtree of order 1 is either a cluster head of order 1 (i.e., $x\in\Phi^{1}_{c}$) or such that $x\in\Phi^{1}_{a}$, and there is a path from $x$ to a cluster head of order 1. See Figure 3.
Figure 3: Left picture: The $f^{1}$-graph generated on the PPP with the $1$-cycles shown in green bold. The difference between this graph and the $f^{0}$-graph is the addition of new edges from a cycle of order zero to the closest point of the nearest cycle of order zero, shown in green.
Right picture: Cluster subtrees of order 1 obtained by deleting the edges of the $1$-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order 1, and the other vertices have directed paths towards the cluster head of order 1 within the cluster.
Figure 4: Estimation of the (logarithm of the) intensity of the exit points at different levels of the algorithm. The initial point process contains approximately 12,000 points. As shown in the plot, the intensity decays by a factor close to $1/3$ at each level, suggesting that the intensity of the exit points at level $k$ is approximately $(1/3)^{k}$.
Order 2
To define the next point-shift $f^{2}$ on the PPP, we need to introduce some notions similar to those defined earlier in Definition 2.3.
Definition 2.5.
Let $\{C^{1}_{i}\}_{i}$ denote the collection of $1$-cycles, and let $\{S^{1}_{i}\}_{i}$ be the collection of all such pairs of points: $S^{1}_{i}$ is the pair of head points belonging to the cycle $C^{1}_{i}$. The sets $\{S^{1}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{1}_{c}$.
Consider the nearest neighbor map, denoted by NN2, on $\{S^{1}_{i}\}_{i}$, which maps each pair $S^{1}_{i}$ to its nearest neighbor in $\{S^{1}_{i}\}_{i}$. The distance between the sets is again the single linkage distance, $\delta$, defined in (1).
The NN2 graph is a directed graph on $\{S^{1}_{i}\}_{i}$. There is an edge from $S^{1}_{i}$ to $\mathrm{NN2}(S^{1}_{i})$, the nearest neighbor of $S^{1}_{i}$ based on the single linkage distance. Let $\mathrm{NN2}(S^{1}_{i})=S^{1}_{j}$. There is a point, denoted $x_{i}$, in $S^{1}_{i}$ and a point, $x_{j}$, in $S^{1}_{j}$ that achieve the minimum in $\delta(S^{1}_{i},S^{1}_{j})$, denoted by $\mathrm{nn2}(x_{i})=x_{j}$. The point $x_{i}$ is called the exit point of order 1 in the pair $S^{1}_{i}$.
Define $f^{2}$ to be the point-shift on $\Phi^{0}$ that coincides with $f^{1}$ everywhere except for the exit points of order 1. For each such point, say $x\in S^{1}$, where $S^{1}$ is the pair where $x$ is the exit point, define $f^{1}(x)=\mathrm{nn2}(x)$. Again, $f^{2}$ defines a point-shift on $\Phi^{0}$. Its graph is $\mathcal{F/F}$, with cycles of length 2 a.s., as there are no ties a.s. The points belonging to these $2$-cycles are exit points of order 1 which form a sub PP of $\Phi^{1}_{c}$ that will be called cluster heads of order 2 and denoted $\Phi_{c}^{2}$. Let $\Phi_{a}^{2}:=\Phi^{0}\setminus\Phi_{c}^{2}$ denote the vertices of the acyclic part of the $f^{2}$ graph. When deleting the $2$-cycles, i.e. the two directed edges, one gets a collection of finite trees which will be called the cluster subtrees of order 2. See Figure 5.
Figure 5: Left picture: The $f^{2}$-graph generated on the PPP with the $2$-cycles shown in red bold. The difference between this graph and the $f^{1}$-graph is the addition of new edges from a cycle of order 1 to the closest point of the nearest cycle of order 1, shown in green. Right picture: Cluster subtrees of order 2 obtained by deleting the edges of the $1$-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order 2, and the other vertices have directed paths towards the cluster head of order 2 within the cluster.
As illustrated in the definitions of the point-shifts $f^{1}$ and $f^{2}$, for each $k\geq 2$, the point-shift $f^{k}$ can be defined using the previous level point-shift, $f^{k-1}$. More generally, $f^{k}$ is defined by induction as follows:
Induction Assumption
The induction assumption is that:
  • $f^{k-1}$ has been constructed as an $\mathcal{F/F}$ point-shift on $\Phi^{0}$ (in the case of PPP, this property is shown in Theorem 2.8);
  • The k1{k-1}-cycles $\{C^{k-1}_{i}\}$ of the connected components are all of length 2 a.s.;
  • The point process $\Phi^{k-1}_{c}$ of cyclic points of the $f^{k-1}$-graph has an intensity bounded above by $\rho/2^{k-1}$.
Induction step
Let $\{S^{k-1}_{i}\}_{i}$ be the collection pairs of points of $C^{k-1}_{i}$. Let NNk be the nearest neighbor map on $\{S^{k-1}_{i}\}_{i}$, which maps each pair $S^{k-1}_{i}$ to its nearest neighbor pair in $\{S^{k-1}_{i}\}_{i}$, again with respect to the single linkage distance, $\delta$. The NNk graph is the directed graph on $\{S^{k-1}_{i}\}_{i}$ with an edge from $S^{k-1}_{i}$ to $\mathrm{NNk}(S^{k-1}_{i})$, the nearest neighbor of $S^{k-1}_{i}$ based on the single linkage distance.
Let $\mathrm{NNk}(S^{k-1}_{i})=S^{k-1}_{j}$. There is a point, denoted $x_{i}$, in $S^{k-1}_{i}$ and a point, $x_{j}$, in $S^{k-1}_{j}$ that achieve the minimum in $\delta(S^{k-1}_{i},S^{k-1}_{j})$, denoted by $\mathrm{nnk}(x_{i})=x_{j}$. The point $x_{i}$ is called the exit point of order k-1 in the pair $S^{k-1}_{i}$.
One defines $f^{k}$ from $f^{k-1}$ by keeping the images of all points the same except for exit points of order k1k-1. The image of such a point say $x\in S^{k-1}_{i}$ define $f^{k}(x)=\mathrm{nnk}(x)$.
Each connected component of the graph of $f^{k}$ on $\Phi^{0}$ is referred to as a cluster of order k. It will be shown in Corollary 4.6 that these connected components are almost surely $\mathcal{F/F}$, meaning that each cluster in the graph of $f^{k}$ is finite and contains a unique cycle. These cycles are of length 2 a.s. and connect two exit points of order k1k-1, denoted $x$ and $y$, and such that $\mathrm{nnk}(x)=y$ and $\mathrm{nnk}(y)=x$. These cycles are called $k$-cycles.
The vertices in the $f^{k}$-graph that belong to $k$-cycles are called cluster heads of order k. The union of cluster heads of order k forms a sub-PP of $\Phi_{c}^{k-1}$, denoted by $\Phi^{k}_{c}$. Let $\Phi^{k}_{a}=\Phi^{0}\setminus\Phi^{k}_{c}$.
If the edges of the $k$-cycles in the $f^{k}$-graph are deleted, one obtains a collection of directed trees referred to as cluster subtrees of order k. Each vertex $x$ in a cluster subtree of order $k$ is either a cluster head of order $k$ (i.e., $x\in\Phi^{k}_{c}$) or such that $x\in\Phi^{k}_{a}$, and there is a path from $x$ to a cluster head of order k. The assumption that the intensity of $\Phi_{k-1}^{c}$ is at most $\rho/2^{k-1}$ implies that the intensity of $\Phi_{k}^{c}$ is at most $\rho/2^{k}$.
Remark 1.
Denote by $\Phi_{e}^{k}$ the point process of exit points of order $k$, and let the intensity of $\Phi_{e}^{0}$ be $\rho/2$. Under the first level of the $\mathrm{CHN}^{2}$ clustering algorithm, each point in $\Phi_{e}^{0}$ is either a cluster head of order one or is connected to a cluster head of order one, and each resulting cluster contains exactly two cluster heads. Furthermore, when two cluster heads of order one belong to the same cluster of order one, one of them becomes an exit point of order one while the other does not. Therefore, the intensity of exit points of order one is less than $\rho/4$. By the same reasoning, the intensity of exit points of order $k$ is less than $\rho/2^{k+1}$. Figure 4 shows simulation estimates of the intensity of exit points at level $k$. As seen in the figure, the intensity at each level is reduced by approximately a factor of three, suggesting that the estimated intensity of exit points at level $k$ is close to $\rho/3^{k+1}$.
2.2 Construction of the Limiting Point-Shift
This subsection discusses the existence and properties of the limiting point-shift and point-shift graph of the $\mathrm{CHN}^{2}$ clustering algorithm.
Definition 2.6.
Consider the sequence $f^{n}$ of the point-shifts constructed in Subsection 2.1 on a Poisson Point Process (PPP). The fact that the intensity of $\Phi^{i}_{c}$ tends to zero as $i$ tends to infinity, as shown in Subsection 2.1, implies that the sequence $\{\Phi_{a}^{k+1}\setminus\Phi_{a}^{k}\}_{k\geq 0}$, where $\Phi_{a}^{0}=\emptyset$, forms a partition of $\Phi^{0}$. This defines a new point-shift, denoted by $f^{\infty}$. The graph of this point-shift, referred to as the $f^{\infty}$-graph and denoted by $G$, is acyclic, due to the construction of the sequence $f^{n}$.
The $f^{\infty}$ point-shift is called the Centroid Hierarchical Nearest Neighbor Point-Shift ($\mathrm{CHN}^{2}PS$), and its associated graph is called the Centroid Hierarchical Nearest Neighbor Eternal Family Forest ($\mathrm{CHN}^{2}$ EFF). When this forest is a tree, it is referred to as the Centroid Hierarchical Nearest Neighbor Eternal Family Tree ($\mathrm{CHN}^{2}$ EFT).
Theorem 2.7 (Construction of the point-shift graph).
Under the Palm probability of $\Phi^{0}$ and rooted at zero, the $\mathrm{CHN}^{2}$ EFF is the local weak limit of the $f^{n}$-graphs on $\Phi^{0}$ as $n\to\infty$. This limit graph is unimodular.
Proof.
This is assessed under the Palm probability of $\Phi^{0}$. For all balls of radius $h$ centered at 0, the restriction of $f^{n}$-graphs to this ball converges a.s. to a limit when $n$ tends to infinity. This is because the intensity of the points of $\Phi^{i}_{c}$ tends to 0 as $i$ tends to infinity. Hence, there is a random $n$ such that the restriction of the graph $G^{n}$ to this ball is fully determined by $f^{n}$ (this random integer is the largest integer $n$ such that the ball in question is empty of points of $\Phi^{n+1}_{c}$). We will denote by $[G,0]$ the limiting random graph, and by $G^{n},0]$ the $f^{n}$-graph under the Palm probability of $\Phi^{0}$. Since for each $n$, $G^{n},0]$ is a unimodular graph (see [haji_mirsadeghi_baccelli2018]) the limit graph $[G,0]$ is also unimodular (see [Baccelli2017]). ∎
Theorem 2.8.
For each $n\in\mathbb{N}$, the connected components of the graph $G^{n},0]$ are all of type $\mathcal{F/F}$.
Proof.
In Section 4, it will be shown that, for all $n$, there is no infinite path in $G^{n},0]$. This fact demonstrates that, for each $n$, all the clusters of $G^{n},0]$ are finite. Note that the cycle indicated in the properties of the $\mathcal{F/F}$ components (see Theorem 3.10 in [Baccelli2017]) in each cluster of $f^{n}$-graph is the cycle of mutual nearest neighbor points, which is a cycle of length two a.s. ∎
Remark 2.
Theorem 2.8 implies that clusters of level $n$ in the $\mathrm{CHN}^{2}$ algorithm are almost surely finite for each $n$.
In the next proposition, it will be shown that all the connected components of the $\mathrm{CHN}^{2}$ EFF share the same distribution.
Proposition 2.9.
Let $G$ be the $\mathrm{CHN}^{2}$ EFF under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}^{d}$, and let $f^{\infty}$ be $\mathrm{CHN}^{2}PS$ on $G$. Then the distribution of $[G,0]$ is identical to the distribution of $[G,v]$, for all $v$ in the support of $\Phi^{0}$.
Proof.
It is known that there exists a directed graph $G^{\prime}$ with vertex set $\Phi^{0}$, that is constructed in a deterministic and isometry-invariant way, which is a.s. a doubly infinite path (see [Matching]). Consequently, a point-shift on $\Phi^{0}$ can be obtained by mapping each vertex to its subsequent vertex following the direction of this path. This point-shift is a bijection. By applying Mecke’s Point Stationary Theorem (see [haji_mirsadeghi_baccelli2018]), it can be concluded that the distribution of $[G,0]$ is identical to the distribution of $[G,v]$. ∎
Corollary 2.10.
The connected components of the $\mathrm{CHN}^{2}$ EFF, $[G,0]$, belong to the $\mathcal{I/I}$ class of the foil classification theorem. In particular, under the PPP, each connected component is a one-ended tree and all of its foils (equivalence classes induced by the point-shift) are infinite a.s.
Proof.
As mentioned previously, at each level $k$, the equality $\Phi^{0}=\Phi^{k}_{c}\cup\Phi^{k}_{a}$ holds, where $\Phi^{k}_{c}$ denotes the cluster heads and $\Phi^{k}_{a}$ represents the acyclic points of order $n$. For a point belonging to $\Phi^{k}_{a}$ at some level $k$, the number of its descendants is finite and does not change thereafter due to the construction of the $\mathrm{CHN}^{2}$ EFF.
Since the intensity of $\Phi^{k}_{c}$ tends to zero as $k$ approaches infinity, the sets $\{\Phi^{k+1}_{a}\setminus\Phi^{k}_{a}\}_{n\geq 0}$, with $\Phi^{0}_{a}=\emptyset$, form a partition of $\Phi^{0}$. Therefore, almost surely, there exists a level $k$ such that zero belongs to $\Phi^{k+1}_{a}\setminus\Phi^{k}_{a}$. Hence, its descendant tree is almost surely finite.
Consequently, the connected component of zero is in the $\mathcal{I/I}$ class. In other words, the cluster containing the origin in the $\mathrm{CHN}^{2}$ EFF graph is a one-ended tree.
Proposition 2.9 implies that all other connected components are also in the $\mathcal{I/I}$ class. ∎
Remark 3 (Evaporation and $\mathcal{I/I}$ Property).
Corollary 2.10 can also be proved using the evaporation property of the $\mathrm{CHN}^{2}$ point-shift, $f^{\infty}$, on the $\mathrm{CHN}^{2}$ EFF, $[G,0]$. Recall that a point-shift $f$ on a connected component $C\subseteq[G,0]$ evaporates $C$ if, almost surely, there is no point $x\in C$ with arbitrarily large backward iterates under $f$, i.e., $f^{-k}(x)=\emptyset$ almost surely for all sufficiently large $k$ (see [haji_mirsadeghi_baccelli2018] for more details). Corollary 24 in [haji_mirsadeghi_baccelli2018] shows that if a point-shift on a point process evaporates a connected component $C$ of its point-shift graph, then $C$ belongs to the $\mathcal{I/I}$ class. Since the intensity of exit points at level $k$ goes to zero as $k\to\infty$, it follows that $f^{\infty}$ evaporates every connected component of $[G,0]$. Hence, all such components belong to the $\mathcal{I/I}$ class.
2.2.1 The Existence of a Last Universal Common Ancestor at infinity.
Considering the genetic data of the species interpretation discussed in the introduction, one can gain clearer insights by examining the properties of the $\mathrm{CHN}^{2}$ EFF graph. Firstly, the $\mathcal{I/I}$ property of the graph shows that each species (point) belongs to an infinite phylogenetic tree. The $\mathcal{I/I}$ property also shows that there exists a LUCA (Last Universal Common Ancestor) ”at infinity”, from which all species in the cluster are descendants. This raises the question of whether there are multiple LUCAs or just one in the clustering of the Poisson point process. This is equivalent to determining the number of connected components that exist in the $\mathrm{CHN}^{2}$ EFF.
Although the number of connected components is generally unknown, in dimension $d=1$, the following theorem shows that the $\mathrm{CHN}^{2}$ EFF has only one connected component.
Theorem 2.11 (Connectivity of the limiting graph in dimension $d=1$).
In dimension $d=1$, under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}$, the limiting $\mathrm{CHN}^{2}$ EFF is almost surely connected.
Proof.
By Corollary 2.7, the limiting $\mathrm{CHN}^{2}$ EFF $[G,0]$ is unimodular (as a local weak limit of the unimodular graphs $G^{n},0]$; see also [Baccelli2017]). By Corollary 2.10, each connected component of $[G,0]$ is infinite almost surely.
Assume for the sake of contradiction that $[G,0]$ is not connected. Because the construction takes place on the line and each component, at every level, connects only to one of its neighbors on the left or on the right, each connected component occupies a consecutive block of vertices along the line (with no interlacing between components). Hence, distinct components occupy disjoint order-convex blocks along $\mathbb{R}$; since every component is infinite, there can be at most two such blocks—one extending to the left and one to the right.
Let $S$ be the set containing the two extremal vertices separating these components—namely, the rightmost vertex of the left component and the leftmost vertex of the right component. These two vertices are the only vertices that have degree one in the graph and can be defined in a covariant way.
Hence, $S$ is a covariant subset of vertices in the sense of Definition 2.7 in [Baccelli2017]. Furthermore, the decomposition of $[G,0]$ into its connected components defines a covariant partition (Definition 2.9 in [Baccelli2017]). However, Lemma 2.11 (No Infinite/Finite Inclusion) in [Baccelli2017] states that, in a unimodular network, a covariant subset cannot have a finite, non-empty intersection with an infinite element of a covariant partition. Here, $S$ intersects each infinite component of the partition in exactly one vertex, violating this property. Equivalently, by Corollary 2.12 in [Baccelli2017], any covariant subset in an infinite unimodular network is almost surely either empty or infinite.Since $S$ contains exactly two vertices, this leads to a contradiction. Therefore, the limiting $\mathrm{CHN}^{2}$ EFF must be connected almost surely, which completes the proof. ∎
The conjecture is that this graph is connected in each dimension $d$ when constructed from a PPP, i.e., there is only one LUCA in the $\mathrm{CHN}^{2}$ EFF phylogenetic tree of a PPP.
2.2.2 The Cardinality of Descendant Trees
The construction of the $\mathrm{CHN}^{2}$ EFF allows us to calculate bounds on the mean number of species within the descendant tree of a typical cluster at a specific level $k$. Note that the direction of the edges in the graph is from child to parent. The $\mathcal{I/I}$ property (of the connected component of zero), ensures a finite number of descendants for each vertex (species), so it enables us to quantify the number of species in the descendant tree.
More precisely, let the intensity of the cluster heads of level $k$, $\Phi_{c}^{k}$, be $\rho_{k}$. Then the mean of the cardinality of the descendant tree of a point belongs to $\Phi_{c}^{k}$, is $1/\rho_{k}$. It is known from Remark 1 that $\rho_{k}$ tend to zero, as $k$ goes to infinity.
Proposition 2.12.
Let $[G,0]$ be the $\mathrm{CHN}^{2}$ EFF under the Palm probability of $\Phi^{0}\subseteq\mathbb{R}^{d}$. Let $N^{0}$ denote the cardinality of the descendant tree of the root 0. Then $\mathbb{E}[N^{0}]=\infty$. In particular, $N^{0}$ has a heavy tail.
Proof.
Since $[G,0]$ is unimodular, the Mass Transport Principle applies (see [Aldous]). Consider the $\mathrm{CHN}^{2}$ point-shift $f$ oriented from each vertex to its parent. Because every component is one-ended, the parent pointers form an infinite ancestral ray almost surely. Applying the MTP to the transport that sends one unit of mass from each vertex to all its ancestors yields $\mathbb{E}[N^{0}]=\infty$. ∎
Remark 4.
The result stated in Proposition 2.12 remains valid for all vertices in the $\mathrm{CHN}^{2}$ EFF. This is due to Proposition 2.9, which ensures that all connected components share the same distribution.
3 Clustroid Hierarchical Nearest Neighbor Chain Algorithm on Cox Point Processes
The aim of this section is twofold. The first one is to show that the algorithms proposed here can be applied to other point processes than Poisson. The second one is to see how the algorithms in question can be used to detect data aggregates when present.
Clustering algorithms are expected to detect aggregates of points in a dataset. Working with a Poisson point process is useful in this context because the PPP can be viewed as a null hypothesis, i.e., a baseline model that has no structural underlying aggregates and where fine random local properties are leveraged to nevertheless recursively cluster data in function of the dataset only. To examine the behavior of the $\mathrm{CHN}^{2}$ algorithm in scenarios where structural aggregates do occur, a natural testbed is that of Cox point processes.
Specifically, the first example (Example 1) is constructed so that each realization contains distinct aggregates of points (see Remark 5), and aims at showing how the algorithm forms clusters when such Cox aggregates are present. As a result, the properties established in Section 2 (such as the finiteness of clusters at each hierarchical level and the weak convergence of the limiting graph,) continue to hold, see Theorem 3.1. Moreover, it is shown in this example how the $\mathrm{CHN}^{2}$ clustering algorithm effectively detects the presence of aggregates. see Subsection 3.1.
Remark 5.
In many clustering algorithms, the term cluster refers to a distinct subgroup of data points that are more similar to one another than to points outside that subgroup. In contrast, in this paper, the word cluster simply denotes a connected component at a given hierarchical level. To avoid confusion between these two usages, the term aggregates of points will be used in this paper to denote data with distinct groups of similar objects.
Example 1 (Cox point process).
A point process $N$ on a state space $\mathcal{S}$ is called a Cox point process (driven by the random measure $M$) if, conditioned on $M$, it is a Poisson point process with intensity measure $M$ (see, e.g., [Bartek, Section 2.3]). Equivalently, $N$ is a Poisson point process whose intensity measure is itself random.
Let $\Xi\subseteq\mathbb{R}^{d}$ be a random closed subset. Define the random measure $M$ on $\mathbb{R}^{d}$ by
$M(B)=\mu(B\cap\Xi),\qquad B\in\mathcal{B}(\mathbb{R}^{d}),$
where $\mu$ denotes the Lebesgue measure on $\mathbb{R}^{d}$. Let $N$ be the Cox point process driven by $M$. Conditioned on the realization of $\Xi$, $N$ is a homogeneous Poisson point process restricted to the random region $\Xi$; in particular, it does not admit any second-order descending chain. If the random set $\Xi$ is stationary in law (for example, generated from a stationary marked point process), then $N$ is also stationary.
Special case. In the simulations of Figures 6(a) and 6(b), the random set $\Xi$ is chosen as a union of random balls, $\Xi=\bigcup_{x\in\xi}B_{r_{x}}(x)$, where $\xi$ is a stationary point process and $\{r_{x}\}_{x\in\xi}$ are i.i.d. random radii. This construction illustrates the behavior of the $\mathrm{CHN}^{2}$ clustering algorithm when applied to Cox point processes exhibiting spatial aggregation.
The following theorem shows that all properties established for the homogeneous PPP extend to this Cox setting.
Theorem 3.1.
Let $N$ be a stationary point process in $\mathbb{R}^{d}$ that almost surely admits no second-order descending chain (Def. 4.1), and assume the same general-position (no ties) conditions as in Section 2. This includes, in particular, any stationary sub–point process of a homogeneous PPP and the Cox point process defined in Example 1.
Then the following properties hold for $\mathrm{CHN}^{2}$ on $N$:
  1. (i)
    For each $n\in\mathbb{N}$, the connected components of the $f^{n}$-graph are almost surely in the $\mathcal{F/F}$ class (finite with a unique cycle).
  2. (ii)
    The sequence of $f^{n}$-graphs converges locally weakly to a unimodular limit graph.
  3. (iii)
    Under the Palm version of $N$, the connected component of the root in the limiting graph belongs to the $\mathcal{I/I}$ class (almost surely one-ended with all foils infinite).
Moreover, if there exists a nontrivial bijective point-shift on $N$ (so that $[G,o]$ and $[G,v]$ have the same distribution for every vertex $v$), then property (iii) holds for all components; that is, every connected component of the limit is $\mathcal{I/I}$ almost surely.
Proof.
(i) Finiteness of the pre-limit graphs. The proof follows by repeating the argument of Theorem 2.8 with $\Phi^{0}$ replaced by $N$. That proof only uses (a) stationarity and (b) the absence of (second-order) descending chains to rule out infinite paths (via the reduction to second-order chains as in Proposition 4.5). Both hypotheses hold here by assumption, hence every $f^{n}$-component is a.s. $\mathcal{F/F}$
(ii) Unimodular local weak limit. The proof follows by the proof of Corollary 2.7: the vanishing intensity of cyclic points across levels implies local stabilization of the $f^{n}$-graphs on bounded windows, yielding a local weak limit $[G,0]$; unimodularity of the limit then follows from unimodularity of the prelimits and closedness of unimodularity under local weak limits (see also [Baccelli2017]).
(iii) One-endedness of the component of the root. This follows from adapting the proof of Corollary 2.10 (or Remark 3): under the Palm version of $N$, evaporation holds because the intensity of exit points tends to zero as the level increases. Hence, the descendant tree of the root is finite while the ancestral ray is infinite. By the foil-classification theorem for unimodular networks ([Baccelli2017, Thm. 3.10]), the rooted component is $\mathcal{I/I}$ almost surely.
Upgrading to all components under a bijective point-shift. When a bijective point-shift exists, the distributions of $[G,o]$ and $[G,v]$ are identical for every vertex $v$ (cf. Proposition 2.9 for the PPP case). The $\mathcal{I/I}$ property then extends from the Palm root to every vertex, so all components are $\mathcal{I/I}$ almost surely. In particular, this condition is satisfied by the stationary Cox point process of Example 1, since the existence of a bijective point-shift for any stationary non-equidistant point process follows from Timár’s theorem [timar2009, Theorem 1]. ∎
(a) A realization of the Cox $P$ (Example 1) with $S=\{3$ points}\} and $r=\{40,20,30\}$; $\sim$2000 points; merges after six steps.
(b) A realization of the Cox $P$ (Example 1) with $S=\{4$ points}\} and $r=\{30,20,30,30\}$; $\sim$2500 points; merges after seven steps.
Figure 6: CHN2 on two Cox realizations.
3.1 Stopping Criteria for Detecting Data Aggregation in $\mathrm{CHN}^{2}$ Clustering
A classical question in hierarchical clustering is that of the identification of aggregates through the clustering process. It is proposed here to use stopping rules purely based on the lengths of edges used in clustering to assess the presence of aggregates. These stopping rules are defined by comparison to the null hypothesis which is that of a Poisson dataset.
To answer this, the mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm are computed and plotted for each data set in Figure 7. In this figure, the left diagram corresponds to the data set of Figure 6(a), while the right one corresponds to the data set of Figure 6(b). In both diagrams, one can observe a distinct ’jump’ in the mean distances, which can be used to determine the stopping level for clustering.
To make this more quantitative, a ratio-based jump detection method can be used. Let $d_{\text{Cox},k}$ and $d_{\text{Poisson},k}$ denote the mean nearest-neighbor distances at level $k$ for the Cox and Poisson datasets, respectively. Consider the ratio:
$R_{k}=\frac{d_{\text{Cox},k}}{d_{\text{Poisson},k}}.$
By tracking the evolution of $R_{k}$ across levels, one can identify aggregation on the basis of a sudden increases in this ratio. Specifically, aggregation is detected at the first level $k^{*}$ where the relative increase in $R_{k}$ exceeds a predefined threshold, i.e.,
$\frac{R_{k}-R_{k-1}}{R_{k-1}}>\tau.$
The results of this method are illustrated in Figure 8, where significant jumps in the mean distance ratios are observed at specific levels. For the dataset in Figure 6(a), the first pronounced jump occurs at level $6$, while for Figure 6(b), it appears at level $5$. This is for a threshold $\tau=0.3$.
Figure 7: Left Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, serving as a null hypothesis and the data set shown in Figure 6(a). At level 6, a pronounced jump is observed in the Cox case, suggesting visually a reasonable stopping criterion. A smaller jump is also visible at level 5, which may indicate partial detection of early-stage clusters.
Right Diagram: Mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm, for a Poisson dataset, and the data set shown in Figure 6(b). The first significant jump is observed at level 5, suggesting visually a suitable stopping level for this aggregated dataset.
Figure 8: Left Diagram: Ratio of mean distances between clusters at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the Cox dataset shown in Figure 6(a). Clustering is detected at level 6 for a threshold $\tau=0.3$.
Right Diagram: Ratio of mean distances at different levels of the $\mathrm{CHN}^{2}$ clustering algorithm for a Poisson dataset and the dataset shown in Figure 6(b). Clustering is detected at level 5 for a threshold $\tau=0.3$.
4 Proof of $\mathcal{F/F}$ property of the pre-limits graphs of $\mathrm{CHN}^{2}$
In Section 2.1, we claimed that, for every $k\in\mathbb{N}$, the point-shift graph, $f^{k}$, contains no infinite path (component). This section proves this claim. For simplicity, it is first shown that the result holds for $k=1$. Then, it will be shown that the same result holds for all $k$. Before this, Second-order descending chains are introduced and it is shown in Proposition 4.2 that such chains do not exist in the Poisson point process and in Proposition 4.5 that an infinite path in the $f^{1}$-graph is a second-order descending chain.
Definition 4.1.
Let $N$ be a point process. Let $\{x_{n}\}_{n\in\mathbb{N}}$ be an infinite sequence of points in $N$ without repetition. This sequence is called a second-order descending chain if, for all $i\geq 2$, the inequality $d_{i}<\max(d_{i-1},d_{i-2})$ holds, where $d_{i}=\|x_{i+1}-x_{i}\|$.
Remark 6.
A descending chain $\{x_{n}\}_{n\in\mathbb{N}}$ in a point process $N$ is defined by the property that for all $i\geq 1$, the inequality $d_{i}<d_{i-1}$ holds, where $d_{i}=|x_{i+1}-x_{i}|$ (see [Meester] and [Daley]). Any descending chain in $N$ is a second-order descending chain; however, the converse does not hold.
Proposition 4.2.
A $d$-dimensional Poisson point process, $\Phi$, contains no second-order descending chain a.s.
Proof.
By classical arguments, it is enough to show that if $\Phi$ is a homogeneous PPP, the probability of having a second-order descending chain in $\Phi+\delta_{0}$, starting from 0, is zero. Consider the following events:
[…]
and
[…]
where $R$ is an integer. Since $A=\cup_{R=1}^{\infty}A_{R}$, it is enough to show that $\mathbb{P}(A_{R})=0$ for all $R$. For $n\geq 3$, let $A_{R,n}$ be the event of existence of a second-order descending chain, $\{x_{i}\}_{i=0}^{n}$ of length (number of edges) $n$, that is, a chain in $\Phi+\delta_{0}$, with $x_{0}=0$, satisfying the property $d_{i}<\max(d_{i-1},d_{i-2})$ for all $2\leq i\leq n-1$, and such that $d_{0},d_{1}<R$. Then one can write $A_{R}=\cap_{n\geq 3}A_{R,n}$. Since for all $i$, $A_{R,i}\supseteq A_{R,i+1}$, so
$\mathbb{P}(A_{R})=\lim_{n\to\infty}\mathbb{P}(A_{R,n}).$ (2)
Thus it is enough to show that $\lim_{n\to\infty}\mathbb{P}(A_{R,n})=0$.
For $n\geq 3$, let $X_{R,n}$ denote the number of second-order descending chains with length $n$ in $\Phi+\delta_{0}$ with $x_{0}=0$ and with $d_{0},d_{1}<R$. Since $\mathbb{P}(A_{R,n})\leq\mathbb{E}[X_{R,n}]$, it is enough to show that $\lim_{n\to\infty}\mathbb{E}[X_{R,n}]=0$.
For $n\geq 1$, consider the function $g_{n+2}(x,\Phi)$ equal to the number of second-order descending chains with length $n+2$, $x_{0}=0$, and $x_{1}=x\neq 0$, under the assumption that 0 and $x$ are in the support of $\Phi$. Then
$\mathbb{E}[X_{R,n+2}]$ $=\mathbb{E}[\,\sum_{x\in\mbox{supp}(\Phi)\cap B_{R}(0)}g_{n+2}\left(x,\Phi+\delta_{0}\right)]\,$
$=\lambda\int_{B_{R}(0)}\mathbb{E}[\,g_{n+2}\left(x,\Phi+\delta_{0}+\delta_{x}\right)]\,{\mathrm{d}}x,$
where the last equality follows from Campbell’s formula and Slivnyak’s theorem. For all $\Phi$ with a support containing $\{0\}$ and $\{x\}$, consider the function $g^{\prime}_{n+2}(y,\Phi)$ equal to the number of second-order descending chains with length $n+2$, without repeated vertices and such that $x_{0}=0$, $x_{1}=x\neq 0$, and $x_{2}=y$, with $y\neq 0$ and $y\neq x$, when assuming that $y$ is in the support of $\Phi$. By the same arguments as above,
$\mathbb{E}[X_{R,n+2}]$ $=\lambda\int_{B_{R}(0)}\mathbb{E}[\,\sum_{y\in B_{R}(x)\cap\mbox{supp}(\Phi),y\neq 0,y\neq x}g^{\prime}_{n+2}\left(y,\Phi+\delta_{0}+\delta_{x}\right)]\,{\mathrm{d}}x$
$=\lambda\int_{B_{R}(0)}\lambda\int_{B_{R}(x)}\mathbb{E}[\,g^{\prime}_{n+2}\left(y,\Phi+\delta_{0}+\delta_{x}+\delta_{y}\right)]\,{\mathrm{d}}y{\mathrm{d}}x.$
The definition of second-order descending chains, together with the fact that we consider chains without repeated vertices and the stationarity of $\Phi$, imply that the last equation is equal to
$\lambda^{2}\int_{B_{R}(0)}\int_{B_{R}(x)}\mathbb{E}[\,X_{\max(\lVert x\rVert,\lVert y-x\rVert),n}]\,{\mathrm{d}}y{\mathrm{d}}x.$
So the following recursive equation holds for the events $X_{R,n}$:
$\mathbb{E}[X_{R,n+2}]=\lambda^{2}\int_{B_{R}(0)}\int_{B_{R}(x)}\mathbb{E}[\,X_{\max(\lVert x\rVert,\lVert y-x\rVert),n}]\,{\mathrm{d}}y{\mathrm{d}}x.$ (3)
Using this recursive equation, it will be shown in Lemma 4 that the equation (4) holds for $\mathbb{E}[X_{R,n}]$, so when $n$ goes to infinity, the number of descending chain of length $n$ in the Poisson point process goes to zero, and this complete the proof. ∎
Lemma 4.3.
Let $\Phi$ be a $d$-dimensional Poisson point process of intensity $\lambda$, and let $X_{R,n}$ be the number of second-order descending chains of length $n$ with $\max(d_{0},d_{1})<R$. Then for all $n\in\mathbb{N}$,
$\mathbb{E}[X_{R,n}]=\frac{\bigl(\lambda^{2}\,\omega_{d}^{2}\,R^{2d}\bigr)^{\lfloor n/2\rfloor}}{\lfloor n/2\rfloor!}\;\times\;\bigl(\lambda\,\omega_{d}\,R^{d}\bigr)^{\,n\bmod 2},$ (4)
where $\omega_{d}$ is the volume of the unit $d$-dimensional ball.
Proof.
The proof is by induction on $n$. Since the recurrence relation (3) increases $n$ by $2$ at each step, there are two separate induction ”tracks”: one starting from $n=0,2,4,\dots$ and the other starting from $n=1,3,5,\dots$.
For even numbers the base case is $n=0$. The $\mathbb{E}[X_{R,0}]$, is the number of second-order descending chain with length 0 in $\Phi\cup{0}$, with $x_{0}=0$, which is equal to 1, so the base case holds for the even numbers. Assume that (4) holds for an even number $n$. It then follows from (3) that
$\mathbb{E}[X_{R,n+2}]$ $=\lambda^{2}\int_{B_{R}(0)}\int_{B_{R}(x)}\mathbb{E}[\,X_{max(\lVert x\rVert,\lVert y-x\rVert),n}]\,{\mathrm{d}}y{\mathrm{d}}x$
$=\frac{\lambda^{n+2}\omega_{d}^{n}}{n/2!}\int_{B_{R}(0)}\int_{B_{R}(x)}(\max(\lVert x\rVert,\lVert y-x\rVert))^{nd}{\mathrm{d}}y{\mathrm{d}}x.$ (5)
The inner integral is equal to
$\int_{B_{R}(x)}(\max(\lVert x\rVert,\lVert y-x\rVert))^{nd}{\mathrm{d}}y$ $=\int_{\theta\in{\mathbb{S}}^{d-1}}\int_{0}^{R}(\max(\lVert x\rVert,\lVert r\rVert))^{nd}r^{d-1}drd\theta$
$=\int_{\theta\in{\mathbb{S}}^{d-1}}\left(\,\int_{0}^{\lVert x\rVert}\lVert x\rVert^{nd}r^{d-1}dr+\int_{\lVert x\rVert}^{R}r^{(n+1)d-1}dr\right)\,d\theta$
$=\int_{\theta\in{\mathbb{S}}^{d-1}}\left(\,\frac{R^{(n+1)d}}{(n+1)d}+\frac{\lVert x\rVert^{(n+1)d}\times n}{(n+1)d}\right)\,d\theta$
$=\frac{\omega_{d}}{n+1}(\,R^{(n+1)d}+n\lVert x\rVert^{(n+1)d}),$ (6)
where ${\mathbb{S}}^{d}$ denotes the $d$-sphere. Equation (4) together with (4) give
$\mathbb{E}[X_{R,n+2}]$ $=\frac{\lambda^{n+2}\omega_{d}^{n+1}}{(n+1)(n/2!)}\int_{B_{R}(0)}(\,R^{(n+1)d}+n\lVert x\rVert^{(n+1)d}){\mathrm{d}}x$
$=\frac{\lambda^{n+2}\omega_{d}^{n+1}}{(n+1)(n/2!)}(\,\int_{B_{R}(0)}R^{(n+1)d}{\mathrm{d}}x+\int_{B_{R}(0)}n\lVert x\rVert^{(n+1)d}{\mathrm{d}}x)\,$
$=\frac{\lambda^{n+2}\omega_{d}^{n+1}}{(n+1)(n/2!)}(\,\omega_{d}R^{(n+2)d}+\frac{n}{n+2}\omega_{d}R^{(n+2)d})\,$
$=\frac{\lambda^{n+2}\omega_{d}^{n+2}}{\frac{n+2}{2}!}R^{(n+2)d},$
which completes the proof for even numbers.
For odd $n$, the base case is $n=1$. Observe that $\mathbb{E}[X_{R,1}]$, the expected number of second-order descending chains of length $1$, is $\lambda\,\omega_{d}\,R^{d}$. This matches (4) because $\lfloor 1/2\rfloor=0$ and $(1\bmod 2)=1$. Then, by applying the same recurrence relation (3), which again advances indices in steps of two, one proves the result for all odd $n=1,3,5,\dots$. Hence the inductive argument for odd $n$ is identical to the even case, simply starting from $n=1$ instead of $n=0$. This completes the proof.
Definition 4.4.
The definition of the point-shift $f^{1}$ implies that the $f^{1}$-graph, as a directed graph, contains two distinct types of edges on $\Phi_{c}^{0}$. The first type is that of edges between the nearest point w.r.t. $d^{0}$. These edges will be referred to as $d$-type edges of level 0 (the green edges in Figure 9). The other type is that of edges that connect an exit point, $x$, of level zero to ${\mathrm{NN1}}(x)$. These edges are referred to as $\delta$-type edges of level one (the red edges in Figure 9).
Proposition 4.5.
Any infinite path in the $f^{1}$-graph is a second-order descending chain.
Proof.
Assume that the $f^{1}$-graph has an infinite component. Since the $f^{0}$-graph has all its components finite (Proposition 2.1), there must be an infinite path $\mathcal{P}$ on $\{S^{0}_{i}\}_{i}$ (see Definition 2.3), with a structure as depicted in Figure 9.
Figure 9: The structure of an infinite path in $\{S^{0}_{i}\}_{i}$ if it exists. The green arrows show $\mathrm{NN0}$} links ($d$ links of order 0), whereas the red arrows show NN1 links between the exit points of $f^{0}$-cycles and their NN1 points ($\delta$ links of order 1.
More precisely, let $e_{i}$ be the $i$-th edge in $\mathcal{P}$. Then $e_{i}$ can be either of $d$ or $\delta$ type. If $e_{i}$ is of the $d$-type, then, there are two possibilities for $e_{i-1}$ and $e_{i-2}$:
  1. 1.
    $e_{i-1}$ and $e_{i-2}$ are both $\delta$-type edges (see Figure 10.a). In this case, since $\delta$-type edges are consecutive edges in the descending chain on $\{S^{0}_{i}\}_{i}$, one has $\|e_{i-1}\|<\|e_{i-2}\|$. And since $e_{i}$ is a MNN0 edge, then $\|e_{i}\|<\|e_{i-1}\|$. Thus $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)$.
  2. 2.
    $e_{i-1}$ is a $\delta$-type edge and $e_{i-2}$ is a $d$-type edge (see Figure 10.b). Since $e_{i-2}$ and $e_{i}$ are both edges between MNN0 points, this gives
    $\|e_{i-1}\|>\|e_{i}\|\ \mbox{and }\|e_{i-1}\|>\|e_{i-2}\|,$ (7)
    which implies that $\|e_{i}\|<\|e_{i-1}\|=\max(\|e_{i-1}\|,\|e_{i-2}\|)$.
    Figure 10: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on the fact that $e_{i}$ is a $d^{0}$-type edge in the infinite path of the $f^{1}$-graph, if it exists.
If $e_{i}$ is a $\delta$-type edge, then there are three possibilities for $e_{i-1}$ and $e_{i-2}$:
  1. 1.
    both are $\delta$-type edges (see Figure 11.a). In this case, $\|e_{i}\|<\max(\|e_{i-1}\|,\allowbreak\|e_{i-2}\|)$ since these edges are in a descending chain in $\{S^{0}_{i}\}_{i}$.
  2. 2.
    $e_{i-1}$ is a $d$-type edge and $e_{i-2}$ is a $\delta$-type edge (see Figure 11.b). Then
    $\|e_{i-2}\|>\|e_{i}\|\quad\mbox{and}\quad\|e_{i-2}\|>\|e_{i-1}\|,$ (8)
    where the first inequality follows from the descending chain property and the second one holds since $e_{i-1}$ is a MNN0 edge between points. This implies that $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)=e_{i-2}$.
  3. 3.
    $e_{i-1}$ is a $\delta$-type edge and $e_{i-2}$ is a $d$-type edge (see Figure 11.c). It is then easy to show that, in this case also, the inequality $\|e_{i}\|<\max(\|e_{i-1}\|,\|e_{i-2}\|)$ holds.
This completes the proof of Proposition 4.5.
Figure 11: Different possibilities for $e_{i-1}$ and $e_{i-2}$ conditioned on the fact that $e_{i}$ is a $d$-type edge in the infinite path of the $f^{1}$-graph if it exists.
Propositions 4.2 and 4.5 give
Corollary 4.6.
If $\Phi^{0}$ is a Poisson point process, then the components of $f^{1}$-graph of $\Phi^{0}$ all a.s finite.
So far, it has been shown that there is no infinite path or cluster in the $f^{1}$-graph. The following proposition establishes that this holds true for all $f^{n}$-graphs.
Proposition 4.7.
For all $n\in\mathbb{N}$, any infinite path in $f^{n}$-graph is a second-order descending chain.
Proof.
The proof is by induction. Any infinite path in the $f^{0}$-graph is a descending chain in the PPP. Therefore, the base case of the induction is true. For the induction step, suppose there is no infinite path (cluster) in the $f^{n}$-graph. Then, if an infinite path $P^{n+1}$ exists in the $f^{n+1}$-graph, this path lies on $\{S^{n+1}_{i}\}_{i}$ eventually. Thus, there are two types of edges on $P^{n}$: the first type is that between two $\mathrm{MNNn}$} points, and the second type is that of edges that connect an exit point of level $n$, $x$, to ${\mathrm{NN(n+1)}}(x)$. The construction of the $f^{n+1}$ point-shift shows that $P^{n+1}$ forms a descending chain between the $f^{n}$-cycles $\{S_{n}^{i}\}$. The same proof as in Proposition 4.5 shows that if $P^{n}$ exists, it is a second-order descending chain. ∎
5 Related Work and Open Problems
Related Work
This work introduces a hierarchical clustering scheme on a Poisson point process (PPP) via the $\mathrm{CHN}^{2}$ algorithm, generating a sequence of random graphs whose local weak limit under Palm probability defines the $\mathrm{CHN}^{2}$ spanning forest (EFF). Each connected component of the $\mathrm{CHN}^{2}$ EFF on the PPP is one-ended. We briefly situate this construction among related spanning structures.
Factor graphs on PPP.
Ferrari, Landim and Thorisson [Hermann] gave a deterministic (factor) construction on a $d$-dimensional stationary PPP producing a loopless graph, a tree in dimensions $d=2,3$, with one end. Holroyd and Peres [Matching] further showed that in any dimension a factor of the PPP can be a one-ended tree. A factor here is a graph whose vertex set is the support of the point process and whose edge set is a deterministic, isomorphism-invariant function of the configuration. In this sense, the $\mathrm{CHN}^{2}$ EFF on the PPP is also a factor graph (constructed level by level), and it is a one-ended forest.
The Holroyd–Peres proof proceeds by building a nested, hierarchical sequence of partitions of the PPP with finitely many finite components at each level, such that every level-$n$ component is contained in exactly one level-(n+1)(n\!+\!1) component, yielding a single component in the limit; this structure then induces a one-ended tree [Matching]. While their hierarchy is conceptually close to ours, their argument is analytical and does not prescribe an explicit algorithm for forming the components at each level.
Timar [timar2009] extended existence of tree factors beyond PPP to general stationary non-equidistant point processes.
Minimal Spanning Forests.
Aldous and Steele [steele_aldous1992] introduced the minimal spanning forest (MSF) as an infinite-volume analogue of the minimal spanning tree for random points, and conjectured that the MSF of the PPP is almost surely a single one-ended tree. The number of components and their ends remain subtle in general (sharp results are known in special settings; see, e.g., [alexander1995]). The level-0 graph $G^{0},0]$ produced by $\mathrm{CHN}^{2}$ is a subgraph of the MSF on the PPP; beyond level 0, the greedy merging rule of $\mathrm{CHN}^{2}$ leads to structural differences from the MSF.
Random thinning and hierarchical heads.
A thinning-based hierarchical classifier related to clustering is studied in [zuyev_tchoumatchenko2001]: points are independently marked by a distribution $\pi=\{p_{k}\}$, then thinned to retain level-$k$ heads, and lower-level points are attached within Voronoi cells of their nearest higher-level head. This scheme uses additional external randomness, whereas $\mathrm{CHN}^{2}$ is an intrinsic (deterministic) factor of the underlying point process.
Open Problems
We record several questions motivated by the $\mathrm{CHN}^{2}$ point-shift and its EFF.
  1. 1.
    Tree conjecture on PPP. Is the $\mathrm{CHN}^{2}$ EFF a tree (i.e., acyclic and connected componentwise) on the PPP in every dimension? Current results establish one-endedness of components; ruling out cycles at all scales remains open.
  2. 2.
    Point-map probability for $\mathrm{CHN}^{2}$. For a point-shift $f$ acting on configurations, consider the semigroup of translations induced by f-f on probability measures supported on configurations with a point at the origin. The $f$-probability (point-map probability) is the limit of the orbit of this action applied to the Palm distribution, when it exists (see, e.g., [Pointshift]). Does the point-map probability exist for the $\mathrm{CHN}^{2}$ point-shift on the PPP? If so, can it be computed explicitly?
  3. 3.
    Alternative inter-cluster distances. The current $\mathrm{CHN}^{2}$ hierarchy uses the clustroid distance in its merging rule. Analyze the variants obtained by replacing this distance with, for example, Hausdorff distance or single-linkage distance between clusters. For these variants, determine:
    • finiteness of connected components at each hierarchical level;
    • existence and identification of the local weak limit under Palm probability;
    • foil classification (e.g., $\mathcal{F/F}$ at finite levels and $\mathcal{I/I}$ in the limit) and the number of ends of components.
  4. 4.
    Beyond PPP. On stationary Cox or, more generally, stationary non-equidistant point processes (cf. [timar2009]), to what extent do the PPP results persist?
Acknowledgements
This work was supported by the ERC NEMO grant, under the European Union’s Horizon 2020 research and innovation programme, grant agreement number 788851 to INRIA.